### v1

# Closure for Microscale Fluctuations in Chaotic Mixing Fronts

Abstract
Scalar mixing fronts are formed under the stirring and spreading action of a random velocity field. The prediction of concentration statistics below the velocity length scale is an open problem. Here, we show that when fluid agitation is caused by a smooth chaotic flow with weak persistence, the evolution of scalar variance can be uniquely related to the stirring and spreading properties of the flow and the diffusivity of the scalar. Using the spectral description of microscale scalar fluctuations in presence of a large-scale forcing by Kraichnan [1], we uncover the link between variance production and dissipation in a dispersing front. This provides a universal closure for the scalar fluctuation microscale whenever there exists a clear separation between macroscale dispersive forcing and microscale dissipation. We validate the closure model on high-resolution direct numerical simulation of advection-diffusion in chaotic flows of variable strength and persistence. These findings open a new avenue to predict both conservative and reactive transport in incompletely mixed environments, ranging from porous media flows to engineered flows in microfluidic devices.
preprint: APS/123-QED
I Introduction
Figure 1: Scalar front mixed and dispersed by a smooth periodic chaotic flow (sine flow). Scalar concentrations (colorscale) obey Eq. (7) with $\kappa=5\cdot 10^{-6}$ and $\bm{u}$ defined by Eq.(LABEL:eq:sineflow), with $U=0.2$ ($\gamma\approx\sigma^{2}_{\gamma}\approx 0.05$), and $t=300$. All the important length scales appearing in the front mixing problem are shown at real scale in the plot (from small to large). For the considered scenario, the Batchelor scale is $s_{B}\approx 0.01s_{v}$  (1), the modified Batchelor scale is $s_{B}^{\prime}\approx 0.01s_{v}$ (Eq. (28)), the microscale is $s_{m}\approx 0.02s_{v}$ (Eq. (19)), the velocity scale is $s_{v}\equiv 1$, the injection scale is $s_{i}\approx 2.7s_{v}$ (Eq. (36)) and the dispersive scale $s_{d}\approx 1.41s_{v}$ (Eq. (4)).
Solute mixing by chaotic fluid flows occurs in a range of natural and industrial processes, including static mixers [2], turbulence  [3], or porous media [4]. Away from the source, mixing fronts naturally develop between the diluting and dispersing solute and the surrounding fluid, forming concentration gradients over a wide range of length scales (Fig. 1). Such interfaces are observed, for instance, in the fringe of contaminant plumes in the subsurface [5], in density or thermal currents in lakes or coastal areas, or in river junctions [6]. A universal modelling framework that captures the mixing process in such interfaces is still missing, even in the simple case where the solute is passively transported by the flow and the flow is smooth at the scale of interest. In contrast to rough turbulent flows, smooth chaotic flows are characterized by a minimum velocity lengthscale $s_{v}$ below which velocity gradients are constant (Fig. 1). Such conditions are found for laminar flows at low Reynolds numbers or in turbulent flows at high Schmidt numbers, where dissipative scales are larger than mixing scales. Below $s_{v}$, solute mixing occurs by the simultaneous action of fluid deformation and molecular diffusion. The Batchelor scale [7] is the minimum lengthscale at which scalar fluctuations can persist (Fig. 1), given by the length scale at which fluid deformation balances by molecular diffusion as
$s_{B}=\sqrt{\kappa/\gamma},$ (1)
with $\kappa$ the molecular diffusivity and $\gamma$ the mean stretching rate of the chaotic flow. The value of $\gamma$ is dictated by the accumulation of fluid deformations produced by microscale velocity gradients, scaling as $U/s_{V}$ where $U$ is a characteristic velocity scale. When the characteristic persistence time $T$ of velocity gradients is small compared to the characteristic advection time on the velocity scale $s_{V}/U$, the mean stretching rate is expected [1] to scale as
$\gamma\sim T\,(U/s_{v})^{2}.$ (2)
Flow persistence can be quantified by a dimensionless persistence parameter, or Kubo number [8, 9],
$Ku=\frac{UT}{s_{V}}.$ (3)
For $Ku\ll 1$, the flow has weak persistence while for $Ku\gg 1$, the flow has long time correlations. Eqs.(1) and (2) show that scalar fluctuations are maintained below $s_{v}$ whenever ${\kappa}<U^{2}T$, that is, for weakly diffusing or strongly agitated solutes. In unbounded domains, chaotic flows favour the dispersive spreading of solutes leading to the formation of scalar gradients at increasingly large scales
$s_{d}\sim\sqrt{Dt}$ (4)
where $D$ is the dispersion coefficient and $t$ is the time (Fig. 1). Owing to the Taylor-Green-Kubo relation between velocity fluctuations and dispersion, $D$ generally scales as
$D\sim U^{2}T.$ (5)
Since $s_{d}$ grows in time and $s_{v}$ and $s_{B}$ are constant, we have the general lengthscale separation
$s_{d}\gg s_{v}>s_{B}$ (6)
at least asymptotically in time. Practically, resolving scales smaller than $s_{d}$ or $s_{v}$ is often impossible numerically at small $\kappa$, due to the large number of grid cells necessary to resolve $s_{B}$. Thus, one aims at finding macroscale equations that correctly capture small-scale fluctuations into an effective model. In the majority of chaotic flows, such effective description remains largely unknown.
At the microscale, the lamellar theory [10, 11] accurately predicts the evolution of elementary solute filaments, which elongate and dilute by the coupled action of flow stretching and molecular diffusion. In this one-dimensional representation, the filaments are considered to be independent of each other. Thus, the theory does not allow prediction of mixing statistics in the late time when filaments coalesce and where the concentration field is not Gaussian. Aggregation theories, whether considered above[12] or below [13] the velocity length scale $s_{v}$, have attempted to describe the late-time evolution of the scalar statistics through the evolution of the number of filament coalescence. Well constrained in confined systems [14], aggregation theories are difficult to apply in dispersing interfaces because the density of lamella aggregation becomes a function of both space and time  [15]. For chaotic flows in particular, the exponential increase of material lengths in slowly increasing dispersive volumes (Eq.(4)) questions the applicability of a uniform aggregation process [16], since it would lead to the removal of all scalar fluctuations inside the dispersion zone at a finite time. In contrast, we expect chaotic agitation to maintain fluctuations in proportion to the mean concentration gradients (Fig. 1). In particular, numerical simulations Tsang et al. [17], Haynes and Vanneste [18] in periodic domains have shown that the decay rate of scalar variance is controlled by microscale dissipation processes in domains smaller than a few $s_{v}$, while it is controlled by large-scale gradients on larger domains.
Batchelor [7], Kraichnan [1] and later Kalda [19] proposed an alternative description of mixing by smooth chaotic flow, recognising that concentration gradients are transported to smaller scales by random fluid stirring, while molecular diffusion is only effective at small scales to reduce scalar variance. These authors describe this process through an evolution equation for the power density spectrum (PSD) of scalar fluctuations as a function of time and spatial frequency. While relying on a similar description of the fluid deformation process as the lamellar representation, the spectral description is global in that it is not limited to isolated Gaussian lamellae. Thus, it is well adapted to the case of scalar mixing fronts of arbitrary shape, where single lamellar elements are not distinguishable. To the author’s knowledge, a comparison of such a spectral description with direct numerical simulations of scalar mixing in smooth chaotic flows has never been attempted.
While appealing, the spectral description of mixing does not directly inform on the concentration statistics of the mixture, an information which is critical to many transport applications, e.g. for contaminant spreading or reactive transport. Furthermore, there is currently no consensus on the asymptotic shape of the scalar Probability Density Function (pdf) in smooth chaotic flows, with experimental studies pointing to Gamma [20], Exponential [4], Power-law [17, 21] or even Gaussian pdfs [22]. Theoretical results on the advection-diffusion equation [23] constrain the possible shapes of the scalar pdf, but leave open the underlying physical mechanisms. In particular, the relation between the asymptotic scalar pdf moments of a mixing front and the characteristics of the microscale stirring process has not been not established. In this study, we show that scalar fluctuations within a mixing front are controlled both by the stretching properties of the microscale velocity field and by the dispersive spreading of the scalar. Using Kraichnan [1] model, we propose a physical-based closure for microscale fluctuation and validate it against direct numerical simulations of chaotically agitated fronts.
The paper is organised as follows. In section 2, we recall the theoretical foundations allowing to predict the evolution of scalar variance in mixing fronts and the spectrum of microscale scalar fluctuations. Based on these two elements, we propose a theoretical closure for the scalar dissipation. In section 3, we compare these analytical predictions with direct numerical simulations. We consider a smooth and single scale flow field, the random sine flow, as a prototype of chaotic flow for which the persistence, the dispersivity and the stretching rate can be varied. Sine flows have been extensively used to uncover the key feature of chaotic mixing [24, 25]. Their relative simplicity makes them computationally tractable even for large domains and times. We numerically study the evolution of the concentration fluctuation spectrum in the case of the forcing of large-scale scalar gradient. The numerical spectrum shows the characteristic $k^{-1}$ scaling between $s_{d}$ and $s_{B}$, and accurately matches the Kraichnan’s prediction. We show that a macroscale transport equation coupled with the physically-based closure for chaotic fluctuations accurately capture the evolution of mixing fronts. We finally conclude on the potential of our results to model other transport processes.
II Theory
In this section, we recall earlier results on the behaviour of concentration fluctuations in chaotic flows, notably the evolution of scalar variance and the spectrum of scalar fluctuations. Based on these two elements, we then propose a closure model for microscale fluctuations in the presence of a mean scalar gradient.
II.1 Background
II.1.1 Concentration microscale
The transport of a scalar concentration field $c(\bm{x},t)$ under the action of an incompressible velocity field $\bm{u}(\bm{x},t)$ is governed by the advection diffusion equation
$\partial_{t}c+\bm{u}\cdot\nabla c=\kappa\nabla^{2}c.$ (7)
Taking a Reynolds decomposition of the concentration field between ensemble mean and fluctuations reads
$c(\bm{x},t)$ $=$ $\bar{c}(\bm{x},t)+c^{\prime}(\bm{x},t),$ (8)
$\bm{u}(\bm{x},t)$ $=$ $\bar{\bm{u}}(\bm{x},t)+\bm{u}^{\prime}(\bm{x},t),$ (9)
where $\bar{\bullet}$ is the ensemble mean of the microscale chaotic agitation process, of length-scale $\leq s_{v}$. $\bar{\bullet}$ can also be interpreted as the result of a low-pass filtering with characteristic scale $s_{v}$.
Due to this filtering, we may assume a separation of length and times scales between the mean field $\bar{c}(\mathbf{x})$ and the fluctuation $c^{\prime}(\mathbf{x},t)$, such that the fluctuation satisfies the inhomogeneous ADE
$\partial_{t}c^{\prime}+\bm{u}\cdot\nabla c^{\prime}-\kappa\nabla^{2}c^{\prime}=f(\mathbf{x}),\quad f(\mathbf{x})\equiv-\bm{u}\cdot\nabla\bar{c}-\kappa\nabla^{2}\bar{c},$ (10)
where the source $f$ has zero mean. For temporally-periodic flows with flow period $T$, the fluctuation $c^{\prime}$ in a zero mean field ($\bar{c}=f=0$) is comprised of a finite set of $N$ strange eigemmodes [26, 21, 27], which are Floquet modes of the AD operator $\mathcal{L}$ of the ADE (7) that manifest as exponentially decaying patterns that control the mixing process:
$c^{\prime}(\mathbf{x},t)=\sum_{k=0}^{K}\alpha_{k}(t)\varphi_{k}(\mathbf{x},t)e^{-\lambda_{k}t},$ (11)
where the (possibly complex) eigenmode $\varphi_{k}(\mathbf{x},t)$ is $T$-periodic in time, and the eigenvalue $\lambda_{k}$ (with $|\lambda_{k}|\leqslant|\lambda_{k+1}|$) controls the persistence of the mode. The coefficient $\alpha_{k}$ is governed by the initial condition $c^{\prime}(\mathbf{x},0)$ as is time-dependent due to the non-self adjoint nature of $\mathcal{L}$, where spectral energy is transferred toward the lowest energy (low $k$) modes with time [28]. With time, only the dominant non-trivial mode $\varphi_{0}$ persists, leading to the observation [21, 26] of convergence of $c^{\prime}$ to a persistent yet decaying spatio-temporal pattern in the case of time-period flows. Note strange eigenmode analogues also exist for aperiodic and open flows [29]. In the presence of a source term such as that given by the mean field in (10), the initial conditions for $c^{\prime}$ decay and the fluctuation converges [30] to
[…]
where $p_{k}(t)$ is also $T$-periodic and typically $|p_{k}|\geq|p_{k+1}|$, hence the low-$k$ modes dominate the fluctuation. Hence, the fluctuation $c^{\prime}$ admits truly persistent spatio-temporal patterns due to continual excitation of the natural mixing modes $\varphi_{k}$ by the mean field $\bar{c}$. Equation (12) shows how the transfer of scalar energy between the mean field and small-scale fluctuations shapes the asymptotic structure of the fluctuation and how the natural mixing modes act to dissipate this scalar energy. Although Eqs. (11)–(12) refer to the concentration field in the real space, we will see below that they have direct analogues in the spectral space. The fluctuations $c^{\prime}$ have zero mean and variance
$\sigma_{c}^{2}=\overline{c^{\prime}c^{\prime}},$ (13)
and the mean concentration obeys the isotropic advection dispersion equation at macroscale [31],
$\frac{\partial\bar{c}}{\partial t}+\bar{\bm{u}}\nabla\bar{c}=\nabla\cdot((\kappa+\bm{D})\nabla\bar{c}),$ (14)
with $\bar{\bm{u}}$ the mean flow velocity over the microscale agitation and $\bm{D}$ is the dispersion coefficient obtained with the assumption of a Fickian macroscale flux expression [32]
$\overline{\bm{u}^{\prime}c^{\prime}}\equiv\bm{D}\nabla\bar{c}.$ (15)
When microscale fluctuations are isotropic, and so may be expressed as $\bm{D}=D\mathbf{I}$ The dispersive limit is reached for any velocity fluctuations that have a finite correlation time and integrable statistics [31, 33]. Multiplying Eq. 7 by $c$ and averaging, one obtains an evolution equation for the variance of concentration fluctuations [33, 32]
$\frac{\partial\sigma_{c}^{2}}{\partial t}+\bar{\bm{u}}\nabla\sigma^{2}_{c}-D\nabla^{2}\sigma^{2}_{c}=-2\kappa\overline{(\nabla c^{\prime})^{2}}+2D(\nabla\bar{c})^{2}.$ (16)
The left hand side is the transport of scalar variance through advective and dispersive fluxes while the right-hand side is the dissipation and production of scalar variance due to molecular diffusion and dispersive motion respectively. As shown by Eq.(16), the production of scalar variance is caused by large-scale dispersive fluxes arising from gradients in the mean concentration $\bar{c}$, which bring unmixed fluid patches in contact with each other. In contrast, the destruction of the scalar variance results from microscale diffusive fluxes. Dimensional arguments imply that dissipation can be expressed as
$\overline{(\nabla c^{\prime})^{2}}\sim\sigma^{2}_{c}/s_{m}^{2},$ (17)
where $s_{m}$ is a concentration microscale [34] characteristic of the microscale mixing process. Providing an expression for $s_{m}$ is the main closure problem. Two limiting scenarios can be distinguished depending on the presence of a large-scale forcing process. The first is the decaying scenario where the scalar homogeneizes according to (7), in the absence of any source or boundary conditions. In that case, there is no production of scalar variance and assuming negligible transport,
$\sigma^{2}_{c}\sim\exp{(-2\kappa\,t/s_{m}^{2})},$ (18)
i.e., an exponential decay of scalar variance.The second case arises in the presence of a mean scalar gradient, for instance in a mixing front which is slowly evolving compared to the microscale agitation process. In this case, the left-hand side of Eq. (16) can be neglected, and there is equilibrium between dissipation and production terms. The scalar variance reaches a steady-state with
$\sigma^{2}_{c}=\frac{D}{\kappa}s_{m}^{2}(\nabla\bar{c})^{2}.$ (19)
Large scale concentration gradients spontaneously develop in unbounded mixing domains, due to dispersive effects (Eq. (42)). Thus, the second scenario is the most general and so is considered herein. Since the minimum scale of fluctuations in chaotic flows is fixed on the Batchelor scale, we may first argue that
$s_{m}\sim s_{B}.$ (20)
Inserting this closure in Eq. (19) leads to $\sigma^{2}_{c}=D(\nabla\bar{c})^{2}/\gamma$, independent of $\kappa$. In contrast, numerical simulations [18] generally show a weak dependence on $\kappa$, suggesting that a closure based on the Batchelor scale only is not satisfactory. In the following, we show instead that the transfer of energy from large to small scales globally controls the amount of scalar variance residing in mixing fronts.
II.1.2 Microscale power spectra
As noted by Batchelor [7], the stretching action of the velocity field transfers scalar fluctuations from larger length-scales toward $s_{B}$. The energy spectrum of concentration fluctuations is $E_{k}(k)=|\tilde{c}(k)|^{2}$, where the tilde symbol stands for the spatial Fourier transform
$\tilde{c}=\int c^{\prime}(\bm{x})e^{i\bm{k}\cdot\bm{x}}\text{d}\bm{x}$ (21)
with $k=|\bm{k}|=2\pi/\lambda$ the wavenumber, inversely proportional to the wavelength $\lambda$. The scalar spectrum was shown to present a distinct shape in the two mixing scenarios considered above. In the case of unforced scalar decay, the spectrum is essentially flat above the dissipation scale [17]. In the case of forced mixing (for example, in the presence of a large-scale concentration gradient), the energy spectrum shows a characteristic $E_{k}\sim k^{-1}$ scaling below the injection scale (for $k>2\pi/s_{v}$), suggesting the transfer of large-scale fluctuations to small scales by the stirring action of the flow [7]. Above the injection scale, the spectrum must be essentially Gaussian, reflecting dispersive processes (Eq. (14)). Intermediate cases with $k$ exponents between 0 and -1 were found in numerical simulations with periodic domains with periodicity comparable with $s_{v}$[17, 35] suggesting the transition between micro- and macro-scale mixing mechanisms. The evolution of the scalar spectrum of $c^{\prime}$ was related to the flow stretching statistics by Kraichnan [1] and independently by Kalda [19] in the case of a $\delta$-time correlated random velocity field. The theory rests on the fact that, in the absence of diffusion, the scalar gradients follow the same evolution as the elongation of material elements $\rho=\ell/\ell_{0}$ (or material surfaces in 3D), with $\ell$ and $\ell_{0}$ the final and initial length (or area in 3D) of a fluid element, respectively, which increases due to fluid deformation. The theory is based on three main elements. The first element is that in a $\delta$-correlated flow, the evolution equation for the pdf of elongation $\Lambda(\ln\rho)$ is
$\frac{\partial\Lambda}{\partial t}+\gamma\frac{\partial\Lambda}{\partial\ln\rho}=\sigma^{2}_{\gamma}/2\frac{\partial^{2}\Lambda}{\partial\ln^{2}\rho}.$ (22)
where $\gamma$ is the mean stretching rate and $\sigma^{2}_{\gamma}$ the variance of the stretching rate [19, 36] defined as
$\gamma$ $=$ $\lim_{t\to\infty}\frac{1}{t}\overline{\ln\rho(t)},$ (23)
$\sigma^{2}_{\gamma}$ $=$ $\lim_{t\to\infty}\frac{1}{t}\overline{(\ln\rho(t)-\gamma)^{2}}.$ (24)
Eq. (22) shows that $\Lambda$ evolves according to a mean drift with the mean stretching rate while broadening due to fluctuation in stretching. Note that in $\delta-$correlated chaotic flows, the stretching variability is related to the dimension of the space $d=2,3,\cdots$ by  [1, 25]
$d=2\gamma/\sigma^{2}_{\gamma}.$ (25)
The second element is that an increase in $\ln\rho$ induces the same increment of wavenumber $\ln k$ in the spectral concentration gradient domain, because fluid stretching is compensated for by transverse compression, which transfers scalar gradients to smaller wavelengths. Thus, in absence of molecular diffusion, the power spectra density of scalar gradients $\Phi(k)\sim E_{k}k$ must follow Eq. (22), with $\ln\rho$ replaced by $\ln k$.
The last element is that molecular diffusion becomes efficient at large $k$ to dissipate energy at a rate $2\kappa k^{2}$, which can be included in the evolution equation as
$\frac{\partial\Phi}{\partial t}+\gamma\frac{\partial\Phi}{\partial\ln k}=\frac{\sigma^{2}_{\gamma}}{2}\frac{\partial^{2}\Phi}{\partial\ln^{2}k}-2\kappa k^{2}\Phi.$ (26)
In essence, such model resembles the lamellar description of mixing [37] with the difference that diffusion now operates on an arbitrary concentration field rather than single Gaussian lamellae. This equation has been studied by Kraichnan [1] and Kalda [19]. In the case of a source of large-scale concentration gradients, the boundary condition is $\Phi(0,t)\equiv\chi_{0}/\gamma$, where $\chi_{0}$ is the rate at which scalar variance is injected into the system from these large-scale gradients and $\gamma$ is the mean stretching rate of an infinitesimal line element [7]. Kalda [19] provides an explicit solution to Eq. (26) in terms of Hankel functions. Kraichnan [1] proposes a simpler equivalent form (5.14 in the original paper)
$E_{k}(k)=\frac{\chi_{0}}{\gamma k}f(2ks_{B}^{\prime})\exp(-2ks_{B}^{\prime}).$ (27)
where the length scale
$s_{B}^{\prime}=\sqrt{\kappa/\sigma^{2}_{\gamma}}.$ (28)
is a modified Batchelor scale, and $f(k)$ is the solution of the Kummer’s equation [1]
$kf^{\prime\prime}-(2k+2\gamma/\sigma^{2}_{\gamma}-1)f^{\prime}+(2\gamma/\sigma^{2}_{\gamma}-1)f=0,\quad f(0)=1.$ (29)
As shown by Kraichnan [1], in two dimensions, $\gamma/\sigma^{2}_{\gamma}=1$, and $f\propto k^{1/2}$ at large $k$.
Equation (27) describes scalar fluctuations in the Batchelor regime, that is, when $\gamma$ is not scale-dependant, and shows that for low wavenumbers the energy decays as $k^{-1}$, with an exponential cutoff at high wavenumbers. In contrast, above the characteristic velocity length scale, the mean stretching rate drops and the flow agitation becomes essentially dispersive (Eq. (16)). In the following, we show that the two descriptions (chaotic and dispersive) can be matched on a particular injection scale $s_{i}$, scale at which the scalar variance is transmitted at higher wavenumbers at a rate $\chi_{0}$ controlled by dispersion.
II.2 Proposed microscale closure
As shown in Eq.(16), scalar fluctuations are produced at large scale by the mean dispersive flux. We define the injection scale $s_{i}$ as the characteristic scale at which scalar fluctuations are produced inside the mixing front. In turn, once produced by dispersive motions, scalar fluctuations are transferred to smaller scales by the cascade process described by Kraichnan [1](Eq. (27)). The amount of scalar variance residing in the system between the injection and the dissipation scale (Eq. (13)) can be obtained by integration, e.g.
$\sigma^{2}_{c}=\frac{1}{\pi}\int_{2\pi/s_{i}}^{\infty}E_{k}(k)\,\text{d}k,$ (30)
where $E_{k}$ is given by Eq. (27). The integral is analytical when $f=1$, and gives
$\sigma^{2}_{c}\approx\frac{\chi_{0}}{\pi\gamma}|Ei(-4\pi s_{B}^{\prime}/s_{i})|,$ (31)
where $Ei$ is the exponential integral. We show in Fig. 6 that Eq. (31) accurately describes numerical simulations, even in the case of two-dimensional flows where $f\neq 1$. This is explained by the primary importance of the exponential cutoff in setting the value of the integral (34), rather than the exact shape at intermediate wavenumbers.
The value of $\chi_{0}$ is controlled by the large-scale dispersion process. At large scale, the variance production rate due to mean scalar gradients (Eq. (16)) is:
$\frac{\text{d}}{\text{d}t}\sigma^{2}_{c}=2D(\nabla\bar{c})^{2}.$ (32)
Assuming that this is entirely transmitted to smaller scales, then
$\chi_{0}\equiv\pi\frac{\text{d}}{\text{d}t}\sigma^{2}_{c}=2\pi D(\nabla\bar{c})^{2}.$ (33)
Combining this argument with Eq. (31) provides a complete closure for microscale scalar variance
$\sigma_{c}^{2}\approx\frac{2D(\nabla\bar{c})^{2}}{\gamma}\left|Ei(-4\pi s_{B}^{\prime}/s_{i})\right|.$ (34)
Eq. (34) indicates that the variance of scalar fluctuations in the presence of a mean scalar gradient weakly depends on the strength of molecular diffusion, since the exponential integral $Ei(x)\sim-\ln x$ when $x<1$.
Alternatively, the closure can be formulated on the concentration microscale (Eq. (19)) is
$s_{m}^{2}\approx 2s_{B}^{2}|Ei(-4\pi s_{B}^{\prime}/s_{i})|.$ (35)
Hence, Eq. (35) appears as a correction factor to the simple anzatz $s_{m}\sim s_{B}$ (Eq. (20)). $s_{m}$ can be larger or lower than $s_{B}$ depending on the ratio of injection versus dissipation scales $s_{i}/s_{B}^{\prime}$. Numerical inspection of Eq. (19) shows that $s_{m}>s_{B}$ when $s_{i}/s_{B}^{\prime}>22.7$, and reversely. Since this ratio can be interpreted as a root-squared Péclet number ($Pe\equiv(s_{B}^{\prime}/s_{i})^{2}$), the critical microscale where $s_{m}\approx s_{B}$ is found around $Pe\approx 515$.
Precising the value of $s_{i}$ is the final ingredient of the microsscale closure problem. We envision $s_{i}$ as the typical scale that divides the dispersive scales above $s_{v}$ from the deformation scales below $s_{v}$. It can be estimated as the scale at which the characteristic time to disperse $(s_{i}^{2}/D)$ is of the order of the characteristic time to deform $(\gamma^{-1})$ :
$s_{i}\sim\sqrt{D/\gamma}\propto s_{V},$ (36)
where we have used Eqs. (2) and (5). Note that $s_{i}$ has the same form as the Batchelor scale (Eq. (1)), but with the diffusion coefficient replaced by the dispersion coefficient. Note that this scale has previously been reported to be important in scalar decay through periodic domains, with mixing rates controlled either by local fluid deformation or by global dispersion, respectively below and above a critical domain size, estimated between $2$ and $3s_{v}$ [17, 38]. Recall finally that the theoretical prediction for the fluctuation variance (or equivalently the fluctuation microscale) relies on the following assumptions: (i) the flow is $\delta$ correlated in time, (ii) $\gamma$ is constant (advection is chaotic and scale independent) and (iii) $s_{B}\ll s_{i}$.
In the following section, we numerically test the accuracy of this prediction in a periodic, two-dimensional, smooth, and chaotic flow, the random sine flow. Transport is solved for a fixed mean gradient in a periodic domain (Sections  III.4 and III.5) and in the case of a dispersing front in an infinite domain (Section III.6).
III Comparison to direct numerical simulations
III.1 Methods
To numerically test Eq.(34), we first consider advection-diffusion of passive scalar in the presence of a fixed large-scale gradient $\nabla\bar{c}\equiv\bm{g}$. As done for Eq.(8), the scalar concentration may be split into mean field and fluctuations
$c(x,t)=\bar{c}+c^{\prime}=\bm{g}\cdot\bm{x}+c^{\prime}(x,t)$ (37)
Inserting (37) into the advection diffusion equation (7) leads to
$\partial_{t}c^{\prime}+\bm{u}\cdot\nabla c^{\prime}=\kappa\nabla^{2}c^{\prime}-\bm{u}\cdot\bm{g}$ (38)
As mentioned above, the presence of a source term in the advection-diffusion equation creates a forcing for scalar fluctuations on the scale of $\bm{u}$. Setting $g=0$ recovers the unforced transport equation (7). Without loss of generality, we focus on a mixing front orientated in the $x$ direction, for which the forced gradient is
$\bm{g}=g\,\bm{e}_{x},$ (39)
thus leading to a source term equal to $-gu(y)\bm{e}_{x}$.
III.2 Velocity field
We consider a generic smooth chaotic velocity field, the two-dimensional Zeldovich sine flow [26], or random wave flow. The sine flow is a continuous transformation operating on a periodic domain $[0,s_{v}]\times[0,s_{v}]$, which is defined as $\bm{u}=(u,v)$, with
$v(x,t)$ $=$ $\left\{\begin{array}[]{ll}U\sin(2\pi x/s_{v}+\phi_{j})&\text{ for}jT<t<(j+1/2)T,\\ 0&\text{ for }(j+1/2)T<t<(j+1)T\\ \end{array}\right.,$
$u(y,t)$ $=$ $\left\{\begin{array}[]{ll}0&\text{ for }jT<t<(j+1/2)T,\\ U\sin(2\pi y/s_{v}+\psi_{j})&\text{ for }(j+1/2)T<t<(j+1)T\\ \end{array}\right.,$
with $T$ the time period of the flow, $j$ the time period number and $U$ a constant velocity magnitude. $\phi_{i},\psi_{i}\in[0,\pi]$ are random phases that change at each time period. The velocity field is divergence free. When $Ku\to 0$ (Eq. (3)), the flow changes rapidly over time, causing minimal fluid deformations, while for $Ku\to\infty$ the persistence of the flow strongly affects the deformation of the fluid elements, which tend to align in the main shear directions. The limit $Ku\to 0$ is the $\delta$-time correlated flow envisioned in Kraichnan [1], where analytical results about exist (Eq. (27)). Without loss of generality, we set $T=1$ and $s_{v}=1$ such that the unique free parameter of the flow is $U=Ku$, which controls both the stretching and dispersion statistics of the flow and its persistence. Particle dispersion in random sine flow is given in the long time limit by [17]
$D\equiv\langle\Delta x^{2}+\Delta y^{2}\rangle/(2T)={U^{2}T}/{16}$ (42)
where $\Delta x$ and $\Delta y$ are the mean squared displacement caused by a single flow period. In the Kraichnan limit, the mean stretching rate is given by [25, 1]
$\gamma={\pi^{2}U^{2}T}/({8\,s_{v}^{2}}),$ (43)
while the variance of stretching rates is equal to the mean, e.g. $\sigma_{\gamma}^{2}=\gamma$. Note that $\gamma\to 0$ as $U\to 0$, e.g. in the $\delta$-correlated flow limit. In the case of persistent flows ($UT/s_{V}\gg 0$), material lines partially align with shear directions and such that stretching is less efficient and $\gamma<{\pi^{2}U^{2}T}/({8s_{v}^{2}})$. In addition, the alignement of all material interfaces on the same directions tends to reduce stretching heterogeneity ($\sigma^{2}_{\lambda}<\lambda$). More details about this persistent regime can be found in Meunier and Villermaux [36]. In most random chaotic flows, the elongation of material lines in sine flows approximately follows a log-normal distribution with parameters $\gamma t$ and $\sigma^{2}_{\gamma}t$ that depend on the amplitude $U$, such that Eq. (22) is valid.
III.3 Numerical scheme and convergence
The advection diffusion equation (10) is solved using spectral decomposition in Fourier space. We have checked that the scalar field remains periodic even for $g>0$, such that Eq. (38) can be solved on the unit square.
Following Meunier and Villermaux [25], unidirectional advection with source is solved exactly with Fourier transform in a single coordinate, while diffusion is solved in the full spectral domain. Denoting $\tilde{c}_{x}$ (respectively $\tilde{c}_{y}$) the Fourier transform of $c$ in direction $x$ (respectively $y$), and $\Delta t\leq T/2$ the numerical time step, the advection part in the first half-time period ($iT<t<(i+1/2)T$),
$\frac{\partial\tilde{c}_{x}}{\partial t}+ik_{x}u(y)\tilde{c}_{x}=\mathcal{F}_{x}(-g\,u(y))=-g\,\delta(k_{x})u(y).$ (44)
This yields
$\tilde{c}_{x}(k_{x},y,t+\Delta t)=\tilde{c}_{x}(k_{x},y,t)\exp(-ik_{x}u(y)\Delta t)-g\delta(k_{x})u(y)\Delta t.$ (45)
Diffusion is solved sequentially in the full frequency domain,
$\tilde{c}(k_{x},k_{y},t+dt)=\tilde{c}_{x}(k_{x},k_{y},t)\exp(-\kappa k^{2}\Delta t).$ (46)
The second half period ($(i+1/2)T<t<(i+1)T$) is solved similarly, this time without source term, with
$\tilde{c}_{y}(x,k_{y},t+dt)=\tilde{c}_{y}(x,k_{y},t)\exp(-ik_{y}v(x)\Delta t)$ (47)
[…]
We run simulations for a time large enough to reach stationarity (in a statistical sense). This is achieved for $t\gamma\gtrapprox 10$ when the scalar variance oscillates around a mean constant value (Figure 2b). Significant fluctuation remains around the mean behaviour, with a given correlation time scale that depends on the inverse of the mean stretching rate $\gamma^{-1}$ (Figure 2a). This can be explained by the time a scalar fluctuation is transported from large to small scales, at the mean rate $\gamma$.
Figure 2: a) Time evolution of the scalar variance b) Time convergence of the mean $<\sigma^{2}_{c}>$ over 100 realizations of the random phases for various $U$ at fixed $s_{B}^{\prime}=5.3\cdot 10^{-3}$.
After this transient, the simulated scalar field reached a statistically steady state, previously reported as strange eigenmode [39] or persistent pattern [21]. In fact, the asymptotic state of the forced problem is not an eigenmode itself but rather is given by a combination of Floquet modes (Eq. (12)) excited by the source term. The spatial structure of this asymptotic mixing state appears different in the presence or absence of large-scale forcing, as shown in Figure 3, with large-scale gradients clearly appearing in the forced case. This suggests a different repartition of the modes energy in freely decaying or forced scalar transport. In turn, the persistence of the flow, characterised by the value of $UT/s_{v}$, does not strongly affect the heterogeneity of the concentration field (Fig. 3), although it clearly impacts the alignment of material lines with respect to the (vertical) velocity field. Statistical averages are then calculated over time periods of $40\gamma$ and presented below.
III.4 Spectrum at equilibrium
Figure 3: Concentration $c^{\prime}$ for various amplitude $U$ for forced mixing $g=1$ (top and scalar decay $g=0$ (bottom). In the scalar decay case, concentrations are rescaled by the standard deviation.
Figure 4: Comparison between power density spectrum obtained via numerical simulation (dots) and Eq. (27) (lines) at fixed Péclet number for $U=0.1,0.2$ and $0.8$ (subplots a.1) and a.2)) and at fixed $U=Ku$ and various Péclet numbers (subplots b.1) and b.2)). The columns show the same spectrum in log-log and log-lin scale. We used $\gamma/\chi_{0}=\pi$ to normalise $E_{k}$. Note that we normalized $k$ by $1/s_{B}^{\prime}$ in the right column to evidence the universal exponential cutoff.
We plot the spectrum of $c^{\prime}$ for various flow persistence and Péclet numbers in Figure 4. The numerical spectrum clearly shows the $k^{-1}$ scaling at large wavenumber (Figs. 4a.1 and b.1) and the exponential cutoff at small wavenumbers wavenumber (Figs. 4a.2 and b.2) . We test the accuracy of the Kraichnan model (Eq. (27)) combined with the proposed injection mechanism (Eq.(33)) in predicting the numerical spectrum. Given the theoretical stretching statistics and dispersivity of the sine flow (Eqs. (43) and (42)), and the proposed relationship (Eq. (33)), the scalar variance injection rate is
${\chi_{0}}=\gamma\,\frac{g^{2}}{\pi}.$ (48)
[…]
In the case of persistent flows, the theory accurately captures the spectrum in the small-wavenumber limit but fails to describe the high-wavenumber tails, which significantly depart from the exponential prediction. Flow persistence ($U>0$) results in a steeper cutoff than predicted theoretically (Fig. 4 a.2). This may be attributed to the alignment of the material lines in the main shear direction, which tends to limit the heterogeneity of the exponential stretching and push towards a Gaussian cutoff of the spectrum [7]. We will see later that this is, however, not critical for the prediction of scalar fluctuations.
III.5 Scalar fluctuations at equilibrium
In Fig. 5, we plot the scalar fluctuation pdf as a function of flow persistence and Péclet number. Due to the large-scale forcing, the pdf of fluctuation is stationary in time. The pdf shows a characteristic Gaussian shape, as previously observed in forced mixing cases [23, 22]. The Gaussian shape is preserved in all flow persistence and Péclet numbers. Note also that the pdf of scalar fluctuation without large-scale forcing ($g=0$) have thicker tails than the Gaussian (Fig. 5a), which resemble a power law. The Gaussianity of scalar fluctuations in mixing fronts is an important feature because it allows one to accurately capture the distribution of fluctuations by the second moments only, that is, the scalar variance.
Figure 5: Pdf of $c^{\prime}$ as a function of $U$ (Pe=$10^{5}$) and Péclet ($U=0.4$). The black dotted line correspond to the case without large scale forcing, with the concentration normalized by $\sigma_{c}$. Pdf are ensemble averaged over 100 realisations.
In Fig. 6, we report the value of the scalar variance in several flow and transport conditions and compare it with the theoretical closure proposed herein (Eq. (6)). A best fit of the closure (38) to the numerical data gives
$s_{i}\approx 2.7s_{v},$ (49)
for all flow persistence (Fig. 6). This value is close to the critical domain size that divides globally controlled from locally controlled variance decay rates in sine flows (between 2 and $3s_{v}$  [17, 38]), confirming the validity of the proposed scaling Eq. (36). Note that the value of $s_{i}$ may differ in other flows than the sine flow.
Even if Eq. (27) relies on the assumption of $\delta$-correlated flows, we found that the prediction for scalar variance is excellent even in the case of correlations. This is due to the importance of the universal range $k^{-1}$ in determining the integral (16) when $s_{B}^{\prime}\ll s_{i}$. The divergence between curves at large $s_{B}^{\prime}$ (small $(s_{i}/s_{B}^{\prime})^{2}$) is explained by the increasing role of the cutoff shape in the integral, which departs from exponential at finite persistence time (Fig. 4).
Figure 6: Dependence of the variance of scalar fluctuations $\sigma^{2}_{c}$ upon modified Batchelor scale $s_{B}^{\prime}$ (Eq. (28)) and flow amplitudes ($U$) in the sine flow (colored lines with markers). The theoretical approximation Eq. (31) is plotted in thick dashed line while the exact numerical integration of Eq. (34) is plotted as a thin continuous line. In the sine flow, $\chi_{0}/\gamma=1/\pi$ (Eq. (48)) The injection scale $s_{i}$ is fixed to $s_{i}=2.7s_{v}$, to match the numerical data.
It is striking that even at large Péclet numbers, microscale fluctuations remain limited compared to macroscale scalar gradients. For example, at $Pe=1000$, scalar fluctuations are expected to be less than 10% of the macroscopic gradient. At $Pe=10^{8}$, the fluctuations hardly reach 60% of the macroscopic gradient.
III.6 Evolution of a mixing front
Now, we test the capacity of the proposed closure Eq. (35) to capture the evolution of a unidirectional mixing front, agitated by a chaotic microscale flow (the sine flow), and that spreads to large scales due to dispersion. Since $\bar{c}$ follows the macrodispersion equation with $\overline{\bm{u}}=0$, its solution for an initial sharp step function is
$\bar{c}=(1+\text{erf}(x/\sqrt{4Dt})/2.$ (50)
The evolution of the local mean concentration is well captured by the dispersive equation (Fig. 7a)
Figure 7: Direct numerical simulation of the evolution the moments of a scalar front mixed by a chaotic velocity field (continuous line). (a) Mean concentration and (b) Variance of the concentration (b). The velocity field is the sine flow with parameters $U=0.8$ and $s_{v}=T=1$. Moments are obtained by solving Eq. (7) with $\kappa=7.7\cdot 10^{-7}$, and ensemble averaging over 20 random phases realizations. The theoretical prediction using the closure Eq. (35) is shown in dashed lines.
Using the closure model (19) based on flow stretching statistics, we can deduce the spatio-temporal evolution of scalar variance. Since
$\frac{\partial\bar{c}}{\partial x}=\frac{1}{\sqrt{4\pi Dt}}\exp(-x^{2}/(4Dt)),$ (51)
we have
$\sigma^{2}_{c}(x,t)=\frac{\exp(-x^{2}/(2Dt))}{2\pi t}\left|Ei(-4\pi s_{B}^{\prime}/s_{i})\right|.$ (52)
Eq. (52) accurately capture the evolution of scalar variance along the front at late time (Fig. 7b). At early times ($t\leq 5$), the prediction underestimates the spatial extent of the variance. This is probably due to the non-negligible transport of scalar variance at early time (the left hand side of Eq. (16)), which redistribute fluctuations.
IV Conclusion
We investigated the statistical behaviour of mixing fronts in smooth chaotic flows in the presence of large-scale scalar gradients.
We use the spectral description of mixing in time-uncorrelated flows of Kraichnan [1] to propose a new closure model relating the stretching statistics (mean and variance of the stretching rates) to the local variability of solute concentrations (the scalar variance), depending on the strength of large-scale scalar gradient and the Péclet number. The validity of the approach relies on three key properties of the flow. The first is the smoothness of the flow field, which ensures that stretching is a continuous, scale-independent process below the velocity-length scale. The second one is the scale separation induced by chaotic advection, which implies that the dissipation and production scales of the scalar variance are well separated, the Batchelor scale being well below the characteristic velocity length scale. The third is the rapid time-decorrelation of flow, which ensures that the evolution of stretching statistics can be envisioned as a Fokker-Planck equation (43). We show that this property is, however, not critical, and that a finite amount of flow persistence does not strongly affect the accuracy of the model. In addition, recent results [40] show that $\rho$ follows an Orstein-Uhlenbeck process for steady and transient flows, reducing to Eq. (22) in the Fickian case.
The variance of scalar fluctuations created by a chaotic flow in a mean scalar gradient is a universal function of a Péclet number (Fig. 6), defined as $Pe=\sigma^{2}_{\gamma}s_{i}^{2}/\kappa$, where $\sigma^{2}_{\gamma}$ is the variance of the stretching rates in the chaotic flow (typically of the same order of magnitude of the Lyapunov exponent $\gamma$), $s_{i}$ is a characteristic scale at which scalar variance is produced (typically of the order of the velocity length scale) and $\kappa$ the molecular diffusivity.
We show that the proposed closure captures direct numerical simulations of advection-diffusion transport in a prototype chaotic flow, the sine flow, and is able to predict the evolution of mean and fluctuation of a solute mixing front. Numerical observation also suggests that local concentration statistics are essentially Gaussian, such that the modelling of the mean and variance of concentrations is sufficient to describe the behaviour of scalar plumes in smooth chaotic flows.
The microscopic closure (35) is expected to apply in any smooth single-scale chaotic flow of low persistence. This is the case for steady laminar flows, for instance in Stokes flow through 3D porous media, where the velocity fluctuations induce chaotic advection in the 2D plane transverse to mean flow direction [41]. In such flows, fluid deformations occur on a time scale that is related to the velocity length scale and magnitude ($T\sim s_{v}/U$), such that Eq. (24) simplifies to $\gamma\sim U/s_{v}$. Thus, flow persistence is of order $Ku\sim 1$. In fact, the microscale $s_{m}$ has already been studied in the context of 3D heterogeneous flows at Darcy scale  Kapoor and Gelhar [42]. However, above pore-scale, advection is in general not chaotic [43], and $s_{m}$ is a non-trivial function of time and space, which strongly dependent on initial conditions. In contrast, below the Darcy scale, advection is chaotic and the microscale is a constant, independent of the solute plume geometry. Thus, the closure (35) is expected to be relevant in the context of up-scaling transport from pore to Darcy scale in porous media.
More work is still required to estimate the value of $s_{i}$ in flows other than the sine flow and to confirm its physical origin and universality (Eq. (36)). Numerical test of the closure in 3D-time dependent chaotic flows is also of direct interest. Because it describes the amount of mixing at microscale, the proposed closure opens a new avenue to model non-conservative transport processes, such as reactive processes. Extension to bimolecular reactive transport system can be obtained by focusing on the concentration statistics of two reactive species, with fluctuation product $\overline{c_{A}^{\prime}c_{B}^{\prime}}$ rather than a single one $\overline{c^{\prime}c^{\prime}}$ as done here [44]. This will be the object of a future study.

### v2

@@
-Closure for Microscale Fluctuations in Chaotic Mixing Fronts
+Mixing Fronts in Chaotic Flows
 Abstract
-Scalar mixing fronts are formed under the stirring and spreading action of a random velocity field.
-The prediction of concentration statistics below the velocity length scale is an open problem.
-Here, we show that when fluid agitation is caused by a smooth chaotic flow with weak persistence, the evolution of scalar variance can be uniquely related to the stirring and spreading properties of the flow and the diffusivity of the scalar.
-Using the spectral description of microscale scalar fluctuations in presence of a large-scale forcing by Kraichnan [1], we uncover the link between variance production and dissipation in a dispersing front.
-This provides a universal closure for the scalar fluctuation microscale whenever there exists a clear separation between macroscale dispersive forcing and microscale dissipation.
-We validate the closure model on high-resolution direct numerical simulation of advection-diffusion in chaotic flows of variable strength and persistence.
-These findings open a new avenue to predict both conservative and reactive transport in incompletely mixed environments, ranging from porous media flows to engineered flows in microfluidic devices.
+Mixing fronts develop at the interface of fluids with different solute concentrations leading to jointly evolving macroscopic and microscopic concentration gradients.
+Macroscopic gradients decay along the front through dispersion driven by microscopic velocity heterogeneity.
+The resulting macroscopic fluctuations are transmitted at the micro-scale through stretching and compression, where they are ultimately dissipated.
+While the elementary mechanisms leading to dispersion on the one hand and scalar mixing on the other hand are well understood, predicting how their coupling governs the evolution of concentration statistics within dispersing fronts remains a challenge.
+Here, we use theoretical derivations and numerical simulations of mixing fronts in chaotic flows to link the evolution of scalar variance to the microscale stirring and macroscale spreading properties of the flow.
+We argue that the transfer of energy between the macroscopic and microscopic scales operates at a characteristic length scale $s_{i}$, that we derive by comparing the characteristic times for scalar variance decay from dispersion and mixing.
+This leads to a closed expression for the concentration variance, which captures the results of numerical simulations with no fitting parameters for a broad range of Peclet numbers.
+For increasing stirring persistence, the numerical simulations deviate from the theory in the small Peclet range, which we explain qualitatively.
+These findings open a new avenue for predicting both conservative and reactive transport in mixing fronts, ranging from porous media flows to engineered flows in microfluidic devices.
 preprint: APS/123-QED
 I Introduction
-Figure 1: Scalar front mixed and dispersed by a smooth periodic chaotic flow (sine flow).
+Solute mixing by chaotic fluid flows occurs in a range of natural and industrial processes, including static mixers [1], turbulence  [2], or porous media [3].
+Away from the solute source, mixing fronts naturally develop at the interface between the diluting and dispersing solute and the surrounding fluid, forming concentration gradients over a wide range of length scales (Fig.
+1).
+Such interfaces are observed, for instance, in the fringe of contaminant plumes in laminar flows through porous media [4], or in turbulent density or thermal currents in lakes, coastal areas, or river junctions [5].
+Describing concentration fluctuations in mixing fronts is of particular importance in the context of mixing-driven reactions in laminar [6] or turbulent chaotic flows [7].
+Turbulent and laminar chaotic flows differ by the presence or absence of wide spectrum of velocity gradient scales.
+Figure 1: Scalar front mixed and dispersed by a single-scale periodic chaotic flow (sine flow).
 Scalar concentrations (colorscale) obey Eq.
 (1) with $\kappa=5\cdot 10^{-6}$ and $\bm{u}$ defined by Eq.(LABEL:eq:sineflow), with $U=0.2$ ($\gamma\approx\sigma^{2}_{\gamma}\approx 0.05$), and $t=300$.
-All the important length scales appearing in the front mixing problem are shown at real scale in the plot (from small to large).
-For the considered scenario, the Batchelor scale is $s_{B}\approx 0.01s_{v}$  (1), the modified Batchelor scale is $s_{B}^{\prime}\approx 0.01s_{v}$ (Eq.
-(28)), the microscale is $s_{m}\approx 0.02s_{v}$ (Eq.
-(19)), the velocity scale is $s_{v}\equiv 1$, the injection scale is $s_{i}\approx 2.7s_{v}$ (Eq.
-(36)) and the dispersive scale $s_{d}\approx 1.41s_{v}$ (Eq.
-(4)).
-Solute mixing by chaotic fluid flows occurs in a range of natural and industrial processes, including static mixers [2], turbulence  [3], or porous media [4].
-Away from the source, mixing fronts naturally develop between the diluting and dispersing solute and the surrounding fluid, forming concentration gradients over a wide range of length scales (Fig.
-1).
-Such interfaces are observed, for instance, in the fringe of contaminant plumes in the subsurface [5], in density or thermal currents in lakes or coastal areas, or in river junctions [6].
-A universal modelling framework that captures the mixing process in such interfaces is still missing, even in the simple case where the solute is passively transported by the flow and the flow is smooth at the scale of interest.
-In contrast to rough turbulent flows, smooth chaotic flows are characterized by a minimum velocity lengthscale $s_{v}$ below which velocity gradients are constant (Fig.
-1).
-Such conditions are found for laminar flows at low Reynolds numbers or in turbulent flows at high Schmidt numbers, where dissipative scales are larger than mixing scales.
-Below $s_{v}$, solute mixing occurs by the simultaneous action of fluid deformation and molecular diffusion.
-The Batchelor scale [7] is the minimum lengthscale at which scalar fluctuations can persist (Fig.
-1), given by the length scale at which fluid deformation balances by molecular diffusion as
-$s_{B}=\sqrt{\kappa/\gamma},$ (1)
-with $\kappa$ the molecular diffusivity and $\gamma$ the mean stretching rate of the chaotic flow.
-The value of $\gamma$ is dictated by the accumulation of fluid deformations produced by microscale velocity gradients, scaling as $U/s_{V}$ where $U$ is a characteristic velocity scale.
-When the characteristic persistence time $T$ of velocity gradients is small compared to the characteristic advection time on the velocity scale $s_{V}/U$, the mean stretching rate is expected [1] to scale as
-$\gamma\sim T\,(U/s_{v})^{2}.$ (2)
-Flow persistence can be quantified by a dimensionless persistence parameter, or Kubo number [8, 9],
-$Ku=\frac{UT}{s_{V}}.$ (3)
-For $Ku\ll 1$, the flow has weak persistence while for $Ku\gg 1$, the flow has long time correlations.
-Eqs.(1) and (2) show that scalar fluctuations are maintained below $s_{v}$ whenever ${\kappa}<U^{2}T$, that is, for weakly diffusing or strongly agitated solutes.
-In unbounded domains, chaotic flows favour the dispersive spreading of solutes leading to the formation of scalar gradients at increasingly large scales
-$s_{d}\sim\sqrt{Dt}$ (4)
-where $D$ is the dispersion coefficient and $t$ is the time (Fig. 1). Owing to the Taylor-Green-Kubo relation between velocity fluctuations and dispersion, $D$ generally scales as
-$D\sim U^{2}T.$ (5)
-Since $s_{d}$ grows in time and $s_{v}$ and $s_{B}$ are constant, we have the general lengthscale separation
-$s_{d}\gg s_{v}>s_{B}$ (6)
-at least asymptotically in time.
-Practically, resolving scales smaller than $s_{d}$ or $s_{v}$ is often impossible numerically at small $\kappa$, due to the large number of grid cells necessary to resolve $s_{B}$.
-Thus, one aims at finding macroscale equations that correctly capture small-scale fluctuations into an effective model.
-In the majority of chaotic flows, such effective description remains largely unknown.
-At the microscale, the lamellar theory [10, 11] accurately predicts the evolution of elementary solute filaments, which elongate and dilute by the coupled action of flow stretching and molecular diffusion.
-In this one-dimensional representation, the filaments are considered to be independent of each other.
-Thus, the theory does not allow prediction of mixing statistics in the late time when filaments coalesce and where the concentration field is not Gaussian.
-Aggregation theories, whether considered above[12] or below [13] the velocity length scale $s_{v}$, have attempted to describe the late-time evolution of the scalar statistics through the evolution of the number of filament coalescence.
-Well constrained in confined systems [14], aggregation theories are difficult to apply in dispersing interfaces because the density of lamella aggregation becomes a function of both space and time  [15].
-For chaotic flows in particular, the exponential increase of material lengths in slowly increasing dispersive volumes (Eq.(4)) questions the applicability of a uniform aggregation process [16], since it would lead to the removal of all scalar fluctuations inside the dispersion zone at a finite time.
-In contrast, we expect chaotic agitation to maintain fluctuations in proportion to the mean concentration gradients (Fig.
-1).
-In particular, numerical simulations Tsang et al.
-[17], Haynes and Vanneste [18] in periodic domains have shown that the decay rate of scalar variance is controlled by microscale dissipation processes in domains smaller than a few $s_{v}$, while it is controlled by large-scale gradients on larger domains.
-Batchelor [7], Kraichnan [1] and later Kalda [19] proposed an alternative description of mixing by smooth chaotic flow, recognising that concentration gradients are transported to smaller scales by random fluid stirring, while molecular diffusion is only effective at small scales to reduce scalar variance.
-These authors describe this process through an evolution equation for the power density spectrum (PSD) of scalar fluctuations as a function of time and spatial frequency.
-While relying on a similar description of the fluid deformation process as the lamellar representation, the spectral description is global in that it is not limited to isolated Gaussian lamellae.
-Thus, it is well adapted to the case of scalar mixing fronts of arbitrary shape, where single lamellar elements are not distinguishable.
-To the author’s knowledge, a comparison of such a spectral description with direct numerical simulations of scalar mixing in smooth chaotic flows has never been attempted.
-While appealing, the spectral description of mixing does not directly inform on the concentration statistics of the mixture, an information which is critical to many transport applications, e.g.
-for contaminant spreading or reactive transport.
-Furthermore, there is currently no consensus on the asymptotic shape of the scalar Probability Density Function (pdf) in smooth chaotic flows, with experimental studies pointing to Gamma [20], Exponential [4], Power-law [17, 21] or even Gaussian pdfs [22].
-Theoretical results on the advection-diffusion equation [23] constrain the possible shapes of the scalar pdf, but leave open the underlying physical mechanisms.
-In particular, the relation between the asymptotic scalar pdf moments of a mixing front and the characteristics of the microscale stirring process has not been not established.
-In this study, we show that scalar fluctuations within a mixing front are controlled both by the stretching properties of the microscale velocity field and by the dispersive spreading of the scalar.
-Using Kraichnan [1] model, we propose a physical-based closure for microscale fluctuation and validate it against direct numerical simulations of chaotically agitated fronts.
+The mixing scales appearing in mixing fronts are shown at real scale in the plot (from small to large).
+For the considered scenario, the Batchelor scale is $s_{B}\approx 0.01s_{v}$, the modified Batchelor scale is $s_{B}^{\prime}\approx 0.01s_{v}$, the velocity scale is $s_{v}\equiv 1$, the injection scale is $s_{i}\approx 2.7s_{v}$ and the dispersive scale $s_{d}\approx 1.41s_{v}$.
+While classical scenarios of forced turbulent mixing in closed domains [8, 9, 10], or solute dilution close to a solute source [11] have been well studied, mixing in dispersing fronts [12] is less understood.
+In scalar turbulence, most existing models have focused on “homogeneously heterogeneous” mixing conditions [8], or micromixing, where the existence of macroscale dispersive transport, or macromixing, is neglected [7].
+In turbulent scalar plumes, Baldyga and Bourne [12] suggest that there exists a competition between macro and micro-mixing phenomena in the presence of large scale gradients.
+In laminar chaotic flows with a single velocity length scale $s_{v}$ (Fig.
+1), scalar mixing is well predicted below $s_{v}$ by the local mechanism of stretching-enhanced diffusion [13].
+At the macroscale (above $s_{v}$), it is also well-known that dispersive transport globally controls scalar variance in bounded domains [14, 15].
+However, how these two processes, operating at different scales, dictate the evolution of concentration statistics in dispersing scalar fronts is not fully understood  [12].
+Although mean concentration fields are generally well described by effective Fickian dispersion models [16], a description of scalar fluctuations generally requires additional closure arguments.
+Micromixing closure arguments can be sought in the lamellar theory [17, 18], which predicts the enhancement of scalar gradients by the stretching of isolated solute filaments and, notably, the mixing time after which scalar dissipation is effective.
+Aggregation theories [19, 20] can be used to describe the evolution of concentration statistics in the regime where solute filaments overlap with each other [21].
+However, the aggregation formalism is difficult to apply to dispersing mixing fronts because aggregation becomes a function of both space and time [22].
+Mixing theories describing the evolution of the scalar power-spectra in response to smooth fluid deformations have shown good accuracy in capturing microscale scalar mixing in bounded domains, in both the case of strong [23] or weak [24] stretching persistence, within the use of only a few numerical constants [10].
+However, they have not yet been applied to the case of freely dispersing fronts.
+In this study, we investigate the relation between macroscale dispersive transport processes above $s_{v}$ and microscale mixing processes below $s_{v}$ in laminar single-scale chaotic flows.
+We use the spectral description of microscale mixing to provide a closure for scalar fluctuations within a dispersing front mixed by a smooth chaotic flow.
+We argue that there exists a fixed injection scale $s_{i}$, at which macroscale scalar fluctuations controlled by dispersive spreading are transferred to the microscale, and dissipated by stretching-enhanced diffusion.
+This scale is set by the balance between macroscale dispersive spreading and microscale chaotic mixing.
+This leads to a prediction of the concentration variance within mixing fronts with no fitting parameters.
 The paper is organised as follows.
-In section 2, we recall the theoretical foundations allowing to predict the evolution of scalar variance in mixing fronts and the spectrum of microscale scalar fluctuations.
+In section 2, we recall the equations governing the conservation of scalar variance in mixing fronts and the spectral evolution of microscale scalar fluctuations.
 Based on these two elements, we propose a theoretical closure for the scalar dissipation.
@@
 Their relative simplicity makes them computationally tractable even for large domains and times.
-We numerically study the evolution of the concentration fluctuation spectrum in the case of the forcing of large-scale scalar gradient.
-The numerical spectrum shows the characteristic $k^{-1}$ scaling between $s_{d}$ and $s_{B}$, and accurately matches the Kraichnan’s prediction.
-We show that a macroscale transport equation coupled with the physically-based closure for chaotic fluctuations accurately capture the evolution of mixing fronts.
-We finally conclude on the potential of our results to model other transport processes.
+We numerically study the evolution of the concentration fluctuation spectrum in the case of the existence of a large-scale scalar gradient.
+We show that a macroscale transport equation coupled with the physically-based closure for chaotic fluctuations accurately captures the evolution of mixing fronts.
+We finally conclude on the potential applicability of our results to model other transport processes.
 II Theory
-In this section, we recall earlier results on the behaviour of concentration fluctuations in chaotic flows, notably the evolution of scalar variance and the spectrum of scalar fluctuations.
+II.1 Background
+In this section, we recall earlier results on the behaviour of scalar concentration mean and fluctuations in chaotic flows, notably the evolution of scalar variance and the spectrum of scalar fluctuations.
 Based on these two elements, we then propose a closure model for microscale fluctuations in the presence of a mean scalar gradient.
-II.1 Background
-II.1.1 Concentration microscale
+II.1.1 Macroscale evolution of concentration mean and variance in a dispersing front
 The transport of a scalar concentration field $c(\bm{x},t)$ under the action of an incompressible velocity field $\bm{u}(\bm{x},t)$ is governed by the advection diffusion equation
@@
 $\bar{\bullet}$ can also be interpreted as the result of a low-pass filtering with characteristic scale $s_{v}$.
-Due to this filtering, we may assume a separation of length and times scales between the mean field $\bar{c}(\mathbf{x})$ and the fluctuation $c^{\prime}(\mathbf{x},t)$, such that the fluctuation satisfies the inhomogeneous ADE
-$\partial_{t}c^{\prime}+\bm{u}\cdot\nabla c^{\prime}-\kappa\nabla^{2}c^{\prime}=f(\mathbf{x}),\quad f(\mathbf{x})\equiv-\bm{u}\cdot\nabla\bar{c}-\kappa\nabla^{2}\bar{c},$ (10)
-where the source $f$ has zero mean.
-For temporally-periodic flows with flow period $T$, the fluctuation $c^{\prime}$ in a zero mean field ($\bar{c}=f=0$) is comprised of a finite set of $N$ strange eigemmodes [26, 21, 27], which are Floquet modes of the AD operator $\mathcal{L}$ of the ADE (7) that manifest as exponentially decaying patterns that control the mixing process:
-$c^{\prime}(\mathbf{x},t)=\sum_{k=0}^{K}\alpha_{k}(t)\varphi_{k}(\mathbf{x},t)e^{-\lambda_{k}t},$ (11)
-where the (possibly complex) eigenmode $\varphi_{k}(\mathbf{x},t)$ is $T$-periodic in time, and the eigenvalue $\lambda_{k}$ (with $|\lambda_{k}|\leqslant|\lambda_{k+1}|$) controls the persistence of the mode.
-The coefficient $\alpha_{k}$ is governed by the initial condition $c^{\prime}(\mathbf{x},0)$ as is time-dependent due to the non-self adjoint nature of $\mathcal{L}$, where spectral energy is transferred toward the lowest energy (low $k$) modes with time [28].
-With time, only the dominant non-trivial mode $\varphi_{0}$ persists, leading to the observation [21, 26] of convergence of $c^{\prime}$ to a persistent yet decaying spatio-temporal pattern in the case of time-period flows.
-Note strange eigenmode analogues also exist for aperiodic and open flows [29].
-In the presence of a source term such as that given by the mean field in (10), the initial conditions for $c^{\prime}$ decay and the fluctuation converges [30] to
-[…]
-where $p_{k}(t)$ is also $T$-periodic and typically $|p_{k}|\geq|p_{k+1}|$, hence the low-$k$ modes dominate the fluctuation.
-Hence, the fluctuation $c^{\prime}$ admits truly persistent spatio-temporal patterns due to continual excitation of the natural mixing modes $\varphi_{k}$ by the mean field $\bar{c}$.
-Equation (12) shows how the transfer of scalar energy between the mean field and small-scale fluctuations shapes the asymptotic structure of the fluctuation and how the natural mixing modes act to dissipate this scalar energy.
-Although Eqs.
-(11)–(12) refer to the concentration field in the real space, we will see below that they have direct analogues in the spectral space.
-The fluctuations $c^{\prime}$ have zero mean and variance
-$\sigma_{c}^{2}=\overline{c^{\prime}c^{\prime}},$ (13)
-and the mean concentration obeys the isotropic advection dispersion equation at macroscale [31],
+Due to this filtering, we may assume a separation of length and times scales between the mean field $\bar{c}(\mathbf{x})$ and the fluctuation $c^{\prime}(\mathbf{x},t)$.
+Inserting (2)-(3) in (1), and averaging gives
 $\frac{\partial\bar{c}}{\partial t}+\bar{\bm{u}}\nabla\bar{c}=\nabla\cdot((\kappa+\bm{D})\nabla\bar{c}),$ (4)
@@
 $\overline{\bm{u}^{\prime}c^{\prime}}\equiv\bm{D}\nabla\bar{c}.$ (5)
-When microscale fluctuations are isotropic, and so may be expressed as $\bm{D}=D\mathbf{I}$ The dispersive limit is reached for any velocity fluctuations that have a finite correlation time and integrable statistics [31, 33].
+When microscale velocity fluctuations are isotropic, the dispersion tensor is also isotropic and expressed as $\bm{D}=D\mathbf{I}$.
+Owing to the Taylor-Green-Kubo relation between velocity fluctuations and dispersion, $D$ generally scales as
+$D\sim U^{2}T.$ (6)
+with $U$ and $T$ a characteristic flow agitation velocity and time.
+The dispersive limit is reached for any chaotic flow of finite correlation time and integrable statistics [16, 28].
+Dispersion leads to the spreading of solute and the formation of scalar gradients at increasingly large scales $s_{d}\sim\sqrt{Dt}$ in mixing fronts (Fig.
+1).
+The fluctuating part of the concentration field satisfy the inhomogeneous ADE
+$\partial_{t}c^{\prime}+\bm{u}\cdot\nabla c^{\prime}-\kappa\nabla^{2}c^{\prime}=f(\mathbf{x}),\quad f(\mathbf{x})\equiv-\bm{u}\cdot\nabla\bar{c}-\kappa\nabla^{2}\bar{c},$ (7)
+where the source $f$ has zero mean.
 Multiplying Eq.
-7 by $c$ and averaging, one obtains an evolution equation for the variance of concentration fluctuations [33, 32]
+7 by $c^{\prime}$ and averaging, one obtains an evolution equation for the variance of concentration fluctuations $\sigma_{c}^{2}=\overline{c^{\prime}c^{\prime}}$  [28, 27]
 $\frac{\partial\sigma_{c}^{2}}{\partial t}+\bar{\bm{u}}\nabla\sigma^{2}_{c}-D\nabla^{2}\sigma^{2}_{c}=-2\kappa\overline{(\nabla c^{\prime})^{2}}+2D(\nabla\bar{c})^{2}.$ (8)
 The left hand side is the transport of scalar variance through advective and dispersive fluxes while the right-hand side is the dissipation and production of scalar variance due to molecular diffusion and dispersive motion respectively.
-As shown by Eq.(16), the production of scalar variance is caused by large-scale dispersive fluxes arising from gradients in the mean concentration $\bar{c}$, which bring unmixed fluid patches in contact with each other.
+As shown by the right-hand side of Eq.(8), the production of scalar variance is caused by large-scale dispersive fluxes arising from gradients in the mean concentration $\bar{c}$, which bring unmixed fluid patches in contact with each other.
 In contrast, the destruction of the scalar variance results from microscale diffusive fluxes.
-Dimensional arguments imply that dissipation can be expressed as
-$\overline{(\nabla c^{\prime})^{2}}\sim\sigma^{2}_{c}/s_{m}^{2},$ (17)
-where $s_{m}$ is a concentration microscale [34] characteristic of the microscale mixing process.
-Providing an expression for $s_{m}$ is the main closure problem.
-Two limiting scenarios can be distinguished depending on the presence of a large-scale forcing process.
-The first is the decaying scenario where the scalar homogeneizes according to (7), in the absence of any source or boundary conditions.
-In that case, there is no production of scalar variance and assuming negligible transport,
-$\sigma^{2}_{c}\sim\exp{(-2\kappa\,t/s_{m}^{2})},$ (18)
-i.e., an exponential decay of scalar variance.The second case arises in the presence of a mean scalar gradient, for instance in a mixing front which is slowly evolving compared to the microscale agitation process.
-In this case, the left-hand side of Eq.
-(16) can be neglected, and there is equilibrium between dissipation and production terms.
-The scalar variance reaches a steady-state with
-$\sigma^{2}_{c}=\frac{D}{\kappa}s_{m}^{2}(\nabla\bar{c})^{2}.$ (19)
-Large scale concentration gradients spontaneously develop in unbounded mixing domains, due to dispersive effects (Eq.
+The objective of the paper is to provide a physically-based closure for microscale scalar dissipation in the case of a freely dispersing front mixed by a smooth chaotic flow.
+II.1.2 Microscale power spectra in a smooth chaotic flow
+As noted by Batchelor [23], the stretching action of a smooth chaotic flow transfers scalar fluctuations from $s_{v}$ towards the so-called Batchelor scale $s_{B}=\sqrt{\kappa/\gamma}$, where $\gamma$ the mean stretching rate of the chaotic flow.
+The value of $\gamma$ is dictated by the accumulation of random fluid deformations produced by microscale velocity gradients, scaling as $U/s_{v}$ where $U$ is a characteristic velocity scale.
+Kraichnan [24] described the evolution of the scalar power spectrum (PSD) in the case of weak chaotic flow and in the absence of macroscopic concentration gradients.
+Flow persistence can be quantified by the Kubo number [29, 30], $Ku={UT}/{s_{v}}$, where $T$ is the characteristic persistence time $T$ of velocity gradients.
+When $Ku\ll 1$, the mean stretching rate is expected to scale as
+$\gamma\sim T\,(U/s_{v})^{2}.$ (9)
+In addition, analytical results can be obtained on the shape of the microscale PSD $E_{k}(k)\equiv|\tilde{c}(k)|^{2}$, where the tilde symbol stands for the spatial Fourier transform
+$\tilde{c}=\int c^{\prime}(\bm{x})e^{i\bm{k}\cdot\bm{x}}\text{d}\bm{x}$ (10)
+with $k\equiv|\bm{k}|=2\pi/s$ the wavenumber, inversely proportional to the wavelength, or scale $s$.
+In the case of a large scale source of scalar fluctuations, Kraichnan [24] showed that the PSD follows
+$E_{k}(k)=\frac{\chi_{0}}{\gamma k}{F}_{\gamma/\sigma^{2}_{\gamma}}(2ks_{B}^{\prime})\exp(-2ks_{B}^{\prime}).$ (11)
+where $s_{B}^{\prime}\equiv\sqrt{\kappa/\sigma^{2}_{\gamma}}$ is a modified Batchelor scale defined with the variance of the stretching rate $\sigma^{2}_{\gamma}$ rather than the mean $\gamma$, and the function ${F}_{\gamma/\sigma^{2}_{\gamma}}(k)$ is the solution of a second-order ODE with parameters $\gamma$ and $\sigma^{2}_{\gamma}$ (see Supp.
+Mat.
+for details and Eq.
+(5.14) in Kraichnan [24]).
+In weakly persistent flows, $\sigma^{2}_{\gamma}$ is related to $\gamma$ through the dimension of the space $d$ where mixing occurs [24, 26]:
+$d=2\gamma/\sigma^{2}_{\gamma}.$ (12)
+In particular, when $d=3$, $\gamma=3\sigma^{2}_{\gamma}/2$ and ${F}_{3/2}=1$.
+In turn, when $d=2$, $\gamma=\sigma^{2}_{\gamma}$, $s_{B}=s_{B}^{\prime}$, and ${F}_{1}\approx 1+O(k^{1/2})$ [24].
+Note that both $\sigma^{2}_{\gamma}$ and $\gamma$ have units of time frequencies.
+Eq.
+(11) shows that the PSD decays as $k^{-1}$ for low wavenumbers, with an exponential cutoff at high wavenumbers.
+Note that the Kraichnan spectrum is comparable to Batchelor’s prediction [23] obtained in the case of a quasi-static (long persistence) straining motion, with the difference that the latter has a steeper cutoff ($E_{k}\sim\exp(-k^{2}s_{B}^{2})$) at high wavenumbers due to the absence of stretching fluctuations.
+The Kraichnan [24] theory assumes that the stretching field is smooth ($\gamma,\sigma^{2}_{\gamma}$ are constants independent of scale), thus only capturing scalar fluctuations at scales smaller than $s_{v}$.
+Hence, it does not describe how these microscale fluctuations interact with macroscopic concentration gradients at large scale ($s_{d}\gg s_{v}$).
+A this scale, the mean stretching rate drops and the flow agitation becomes essentially dispersive (Eq.
 (8)).
-Thus, the second scenario is the most general and so is considered herein.
-Since the minimum scale of fluctuations in chaotic flows is fixed on the Batchelor scale, we may first argue that
-$s_{m}\sim s_{B}.$ (20)
-Inserting this closure in Eq.
-(19) leads to $\sigma^{2}_{c}=D(\nabla\bar{c})^{2}/\gamma$, independent of $\kappa$.
-In contrast, numerical simulations [18] generally show a weak dependence on $\kappa$, suggesting that a closure based on the Batchelor scale only is not satisfactory.
-In the following, we show instead that the transfer of energy from large to small scales globally controls the amount of scalar variance residing in mixing fronts.
-II.1.2 Microscale power spectra
-As noted by Batchelor [7], the stretching action of the velocity field transfers scalar fluctuations from larger length-scales toward $s_{B}$.
-The energy spectrum of concentration fluctuations is $E_{k}(k)=|\tilde{c}(k)|^{2}$, where the tilde symbol stands for the spatial Fourier transform
-$\tilde{c}=\int c^{\prime}(\bm{x})e^{i\bm{k}\cdot\bm{x}}\text{d}\bm{x}$ (21)
-with $k=|\bm{k}|=2\pi/\lambda$ the wavenumber, inversely proportional to the wavelength $\lambda$.
-The scalar spectrum was shown to present a distinct shape in the two mixing scenarios considered above.
-In the case of unforced scalar decay, the spectrum is essentially flat above the dissipation scale [17].
-In the case of forced mixing (for example, in the presence of a large-scale concentration gradient), the energy spectrum shows a characteristic $E_{k}\sim k^{-1}$ scaling below the injection scale (for $k>2\pi/s_{v}$), suggesting the transfer of large-scale fluctuations to small scales by the stirring action of the flow [7].
-Above the injection scale, the spectrum must be essentially Gaussian, reflecting dispersive processes (Eq.
-(14)).
-Intermediate cases with $k$ exponents between 0 and -1 were found in numerical simulations with periodic domains with periodicity comparable with $s_{v}$[17, 35] suggesting the transition between micro- and macro-scale mixing mechanisms.
-The evolution of the scalar spectrum of $c^{\prime}$ was related to the flow stretching statistics by Kraichnan [1] and independently by Kalda [19] in the case of a $\delta$-time correlated random velocity field.
-The theory rests on the fact that, in the absence of diffusion, the scalar gradients follow the same evolution as the elongation of material elements $\rho=\ell/\ell_{0}$ (or material surfaces in 3D), with $\ell$ and $\ell_{0}$ the final and initial length (or area in 3D) of a fluid element, respectively, which increases due to fluid deformation.
-The theory is based on three main elements.
-The first element is that in a $\delta$-correlated flow, the evolution equation for the pdf of elongation $\Lambda(\ln\rho)$ is
-$\frac{\partial\Lambda}{\partial t}+\gamma\frac{\partial\Lambda}{\partial\ln\rho}=\sigma^{2}_{\gamma}/2\frac{\partial^{2}\Lambda}{\partial\ln^{2}\rho}.$ (22)
-where $\gamma$ is the mean stretching rate and $\sigma^{2}_{\gamma}$ the variance of the stretching rate [19, 36] defined as
-$\gamma$ $=$ $\lim_{t\to\infty}\frac{1}{t}\overline{\ln\rho(t)},$ (23)
-$\sigma^{2}_{\gamma}$ $=$ $\lim_{t\to\infty}\frac{1}{t}\overline{(\ln\rho(t)-\gamma)^{2}}.$ (24)
-Eq.
-(22) shows that $\Lambda$ evolves according to a mean drift with the mean stretching rate while broadening due to fluctuation in stretching.
-Note that in $\delta-$correlated chaotic flows, the stretching variability is related to the dimension of the space $d=2,3,\cdots$ by  [1, 25]
-$d=2\gamma/\sigma^{2}_{\gamma}.$ (25)
-The second element is that an increase in $\ln\rho$ induces the same increment of wavenumber $\ln k$ in the spectral concentration gradient domain, because fluid stretching is compensated for by transverse compression, which transfers scalar gradients to smaller wavelengths.
-Thus, in absence of molecular diffusion, the power spectra density of scalar gradients $\Phi(k)\sim E_{k}k$ must follow Eq.
-(22), with $\ln\rho$ replaced by $\ln k$.
-The last element is that molecular diffusion becomes efficient at large $k$ to dissipate energy at a rate $2\kappa k^{2}$, which can be included in the evolution equation as
-$\frac{\partial\Phi}{\partial t}+\gamma\frac{\partial\Phi}{\partial\ln k}=\frac{\sigma^{2}_{\gamma}}{2}\frac{\partial^{2}\Phi}{\partial\ln^{2}k}-2\kappa k^{2}\Phi.$ (26)
-In essence, such model resembles the lamellar description of mixing [37] with the difference that diffusion now operates on an arbitrary concentration field rather than single Gaussian lamellae.
-This equation has been studied by Kraichnan [1] and Kalda [19].
-In the case of a source of large-scale concentration gradients, the boundary condition is $\Phi(0,t)\equiv\chi_{0}/\gamma$, where $\chi_{0}$ is the rate at which scalar variance is injected into the system from these large-scale gradients and $\gamma$ is the mean stretching rate of an infinitesimal line element [7].
-Kalda [19] provides an explicit solution to Eq.
-(26) in terms of Hankel functions.
-Kraichnan [1] proposes a simpler equivalent form (5.14 in the original paper)
-$E_{k}(k)=\frac{\chi_{0}}{\gamma k}f(2ks_{B}^{\prime})\exp(-2ks_{B}^{\prime}).$ (27)
-where the length scale
-$s_{B}^{\prime}=\sqrt{\kappa/\sigma^{2}_{\gamma}}.$ (28)
-is a modified Batchelor scale, and $f(k)$ is the solution of the Kummer’s equation [1]
-$kf^{\prime\prime}-(2k+2\gamma/\sigma^{2}_{\gamma}-1)f^{\prime}+(2\gamma/\sigma^{2}_{\gamma}-1)f=0,\quad f(0)=1.$ (29)
-As shown by Kraichnan [1], in two dimensions, $\gamma/\sigma^{2}_{\gamma}=1$, and $f\propto k^{1/2}$ at large $k$.
-Equation (27) describes scalar fluctuations in the Batchelor regime, that is, when $\gamma$ is not scale-dependant, and shows that for low wavenumbers the energy decays as $k^{-1}$, with an exponential cutoff at high wavenumbers.
-In contrast, above the characteristic velocity length scale, the mean stretching rate drops and the flow agitation becomes essentially dispersive (Eq.
-(16)).
-In the following, we show that the two descriptions (chaotic and dispersive) can be matched on a particular injection scale $s_{i}$, scale at which the scalar variance is transmitted at higher wavenumbers at a rate $\chi_{0}$ controlled by dispersion.
-II.2 Proposed microscale closure
-As shown in Eq.(16), scalar fluctuations are produced at large scale by the mean dispersive flux.
-We define the injection scale $s_{i}$ as the characteristic scale at which scalar fluctuations are produced inside the mixing front.
-In turn, once produced by dispersive motions, scalar fluctuations are transferred to smaller scales by the cascade process described by Kraichnan [1](Eq.
-(27)).
-The amount of scalar variance residing in the system between the injection and the dissipation scale (Eq.
-(13)) can be obtained by integration, e.g.
-$\sigma^{2}_{c}=\frac{1}{\pi}\int_{2\pi/s_{i}}^{\infty}E_{k}(k)\,\text{d}k,$ (30)
-where $E_{k}$ is given by Eq. (27). The integral is analytical when $f=1$, and gives
-$\sigma^{2}_{c}\approx\frac{\chi_{0}}{\pi\gamma}|Ei(-4\pi s_{B}^{\prime}/s_{i})|,$ (31)
-where $Ei$ is the exponential integral.
-We show in Fig.
-6 that Eq.
-(31) accurately describes numerical simulations, even in the case of two-dimensional flows where $f\neq 1$.
-This is explained by the primary importance of the exponential cutoff in setting the value of the integral (34), rather than the exact shape at intermediate wavenumbers.
-The value of $\chi_{0}$ is controlled by the large-scale dispersion process. At large scale, the variance production rate due to mean scalar gradients (Eq. (16)) is:
+The low wavenumber PSD of freely dispersing fronts is thus expected to follow
+$E_{k}(k)\sim\exp(-2Dk^{2}t),$ (13)
+. Since $s_{d}$ grows in time and $s_{v}$ and $s_{B}$ are fixed, dispersing fronts show the characteristic separation of lengthscales
+$s_{d}\gg s_{v}\gg s_{B},$ (14)
+at least asymptotically in time.
+In the following, we investigate how to link the Kraichnan description of small scale concentration fluctuations with the macroscopic evolution of large scale concentration gradients in mixing fronts.
+We show that the descriptions of mixing at macroscale (Eq.
+(8)) and at microscale (Eq.
+(11)) can be matched on a particular injection scale $s_{i}$, at which the macroscale scalar fluctuations are transmitted to higher wavenumbers at a rate $\chi_{0}$ controlled by dispersion.
+II.2 Scalar variance in mixing fronts
+Since scalar fluctuations in a freely dispersing front are produced at large scales by the mean dispersive flux (Eq.(8)), we assume that the rate of variance injection $\chi_{0}$ is controlled by the large-scale dispersion process.
+This hypothesis is similar to the spectral equilibrium hypothesis in turbulence, which assumes that the scalar dissipation rate is set by the inertial-convective range, regardless of the extent of the viscous-diffusive range [7].
+The variance production rate due to mean scalar gradients is:
 $\frac{\text{d}}{\text{d}t}\sigma^{2}_{c}=2D(\nabla\bar{c})^{2}.$ (15)
@@
 $\chi_{0}\equiv\pi\frac{\text{d}}{\text{d}t}\sigma^{2}_{c}=2\pi D(\nabla\bar{c})^{2}.$ (16)
-Combining this argument with Eq. (31) provides a complete closure for microscale scalar variance
-$\sigma_{c}^{2}\approx\frac{2D(\nabla\bar{c})^{2}}{\gamma}\left|Ei(-4\pi s_{B}^{\prime}/s_{i})\right|.$ (34)
-Eq.
-(34) indicates that the variance of scalar fluctuations in the presence of a mean scalar gradient weakly depends on the strength of molecular diffusion, since the exponential integral $Ei(x)\sim-\ln x$ when $x<1$.
-Alternatively, the closure can be formulated on the concentration microscale (Eq. (19)) is
-$s_{m}^{2}\approx 2s_{B}^{2}|Ei(-4\pi s_{B}^{\prime}/s_{i})|.$ (35)
-Hence, Eq.
-(35) appears as a correction factor to the simple anzatz $s_{m}\sim s_{B}$ (Eq.
+At the micro-scale, Kraichnan [24] shows that scalar fluctuations transferred from large scales obey a well-defined spectral structure (Eq.
 (11)).
-$s_{m}$ can be larger or lower than $s_{B}$ depending on the ratio of injection versus dissipation scales $s_{i}/s_{B}^{\prime}$.
-Numerical inspection of Eq.
-(19) shows that $s_{m}>s_{B}$ when $s_{i}/s_{B}^{\prime}>22.7$, and reversely.
-Since this ratio can be interpreted as a root-squared Péclet number ($Pe\equiv(s_{B}^{\prime}/s_{i})^{2}$), the critical microscale where $s_{m}\approx s_{B}$ is found around $Pe\approx 515$.
-Precising the value of $s_{i}$ is the final ingredient of the microsscale closure problem.
-We envision $s_{i}$ as the typical scale that divides the dispersive scales above $s_{v}$ from the deformation scales below $s_{v}$.
-It can be estimated as the scale at which the characteristic time to disperse $(s_{i}^{2}/D)$ is of the order of the characteristic time to deform $(\gamma^{-1})$ :
-$s_{i}\sim\sqrt{D/\gamma}\propto s_{V},$ (36)
-where we have used Eqs.
-(2) and (5).
-Note that $s_{i}$ has the same form as the Batchelor scale (Eq.
-(1)), but with the diffusion coefficient replaced by the dispersion coefficient.
-Note that this scale has previously been reported to be important in scalar decay through periodic domains, with mixing rates controlled either by local fluid deformation or by global dispersion, respectively below and above a critical domain size, estimated between $2$ and $3s_{v}$ [17, 38].
-Recall finally that the theoretical prediction for the fluctuation variance (or equivalently the fluctuation microscale) relies on the following assumptions: (i) the flow is $\delta$ correlated in time, (ii) $\gamma$ is constant (advection is chaotic and scale independent) and (iii) $s_{B}\ll s_{i}$.
-In the following section, we numerically test the accuracy of this prediction in a periodic, two-dimensional, smooth, and chaotic flow, the random sine flow.
-Transport is solved for a fixed mean gradient in a periodic domain (Sections  III.4 and III.5) and in the case of a dispersing front in an infinite domain (Section III.6).
+We argue that, in freely dispersing fronts, the transfer of energy from macro- to micro-scales operates at a specific scale $s_{i}$, named the injection scale, which separates dispersive processes (well above the velocity scale) from fluid deformation processes (well below the velocity scale).
+We define $s_{i}$ as the scale at which the characteristic time to destroy variance by dispersion $t_{D}$ is comparable to the characteristic time to destroy variance by stretching-enhanced mixing $t_{\gamma}$.
+In that sense, it is exactly the scale that separates the globally controlled decay–long wavelength mixing mechanism from the locally controlled decay–short wavelength mixing mechanism described by Haynes and Vanneste [15] and Tsang et al.
+[14].
+Note also that $s_{i}$ could also be associated to the “coarsening scale”[31], below which scalar fluctuations are quickly erased by stretching-enhanced mixing.
+While this scale has been investigated n the context of multiscale turbulent flows [19], the precise relation to the scale defined here for smooth single scale flows remains to be established.
+We estimate the injection scale $s_{i}$ as follows. The characteristic time for variance decay by dispersion in a (closed) domain of length $s_{i}$ is
+$t_{D}^{-1}=2\left(\frac{2\pi}{s_{i}}\right)^{2}D$ (17)
+while the characteristic time for variance decay by stretching-enhanced mixing is [32]
+$t_{\gamma}^{-1}=\frac{\gamma^{2}}{2\sigma^{2}_{\gamma}}.$ (18)
+The first equation can be simply obtained with a Fourier transform of the dispersion equation, while the second equation results from log-normal distribution of elongation  [32, 13, 33, 3].
+Using the proportionality between mean and variance of stretching rates (Eq.
+(12)), Eq.
+(18) simplifies to $t_{\gamma}^{-1}=d\gamma/4$ in $d$-dimensional weakly persistent flow.
+Equating $t_{D}$ and $t_{\gamma}$ sets the injection scale to
+$s_{i}=4\pi\sqrt{\frac{D\sigma^{2}_{\gamma}}{\gamma^{2}}}.$ (19)
+For weakly persistent flows in $d=2$ dimensions, we expect $s_{i}=4\pi\sqrt{D/\gamma}$ and in $d=3$ dimensions, $s_{i}=4\pi\sqrt{2D/(3\gamma)}$.
+Note the similarity of these expressions for $s_{i}$ and the Batchelor scale, $s_{B}=\sqrt{\kappa/\gamma}$, with the molecular diffusivity $\kappa$ replaced by the dispersion coefficient $D$.
+Having defined $\chi_{0}$ and $s_{i}$, the amount of microscale scalar variance in a freely dispersing front is directly obtained by integration of the power density spectrum, e.g.
+$\sigma^{2}_{c^{\prime}}=\frac{1}{\pi}\int_{2\pi/s_{i}}^{\infty}E_{k}(k)\,\text{d}k,$ (20)
+where $E_{k}$ is given by Eq. (11). Assuming ${F}_{\gamma/\sigma_{\gamma}^{2}}\approx 1$, the integral is analytical, and gives
+$\sigma^{2}_{c}\approx\frac{\chi_{0}}{\pi\gamma}|Ei(-4\pi s_{B}^{\prime}/s_{i})|,$ (21)
+where $Ei$ is the exponential integral.
+Note that when integrating Eq.
+(8), we have implicitly assumed that the scalar spectrum is well approximated by the Kraichnan model up to the scale $s_{i}$, given by Eq.
+(19).
+Using the expected scaling for $\gamma$ and $D$ (Eqs.
+(9),(6)) in the definition of $s_{i}$ (Eq (19), we expect $s_{i}\sim s_{v}$.
+Our hypothesis is thus justified by the smoothness of the velocity field up to $s_{v}$ (see Fig.
+2).
+While the scalar spectrum likely deviates from the Kraichnan scale close to $s_{i}$, the comparison of Eq.
+(21) with numerical simulations indicate that this has a limited impact on scalar variance.
+Equations (16), (19) and (21) provide a prediction for the magnitude of scalar variance within mixing fronts, based on four independent physical parameters: the molecular diffusion $\kappa$, the dispersivity $D$, the mean and variance of the stretching rate $\gamma$ and $\sigma^{2}_{\gamma}$.
+Note the absence of any numerical constant that cannot be estimated independently, in contrast to existing theories that describe the power density spectrum in turbulent flows [10].
+In Figure 2, we summarise the definitions of all length scales mentioned in the study.
+Figure 2: Sketch of the scalar spectrum of a freely dispersing front, with the dispersive spectrum at macroscale (Eq.
+(13)) and the Kraichnan spectrum at microscale (Eq.
+(11)).
+$s_{d}$ is the dispersion scale, $s_{i}$ is the injection scale, $s_{v}$ is the velocity scale, $s_{B}^{\prime}$ is the modified Batchelor scale and $s_{B}$ is the Batchelor scale.
+The shaded area determines the variability in concentration fluctuations $\sigma^{2}_{c}$
+The closure (21) can be used in the macroscale variance evolution equation (8) as follows. Upon dimensional arguments, the dissipation term can be written as
+$2\kappa\overline{(\nabla c^{\prime})^{2}}\sim-\sigma^{2}_{c}/\tau_{m},$ (22)
+with $\tau_{m}$ a characteristic decay time of scalar variance given by
+$\tau_{m}=\frac{\sigma^{2}_{c}}{2D\left(\nabla\bar{c}\right)^{2}}=\gamma^{-1}\left|Ei(-4\pi s^{\prime}_{B}/s_{i})\right|.$ (23)
+In the first equality, we have used the assumption of quasi-equilibrium where the left-hand side of Eq.
+(8) is negligible and, in the second equality, we have used the closure Eq.
+(21).
+Note that the ratio $s_{i}/s_{B}^{\prime}$ can be associated to the square root of a Péclet number defined as
+$Pe\equiv\left(\frac{s_{i}}{s_{B}^{\prime}}\right)^{2}=8\pi\frac{\sigma^{2}_{\gamma}}{\gamma}\frac{D}{\kappa},$ (24)
+which compares dispersive to diffusive fluxes.
+It is interesting to note that at high $Pe$ (low $s_{B}^{\prime}/s_{i}$), the exponential integral shows a logarithmic growth $|Ei(1/x)|\sim\ln x$, such that Eq.
+(23) takes a form similar to the /expression of lamellar mixing time in chaotic flows [17], $\tau_{m}\sim\gamma^{-1}\ln Pe_{0}$, with $Pe_{0}=s_{0}^{2}\gamma/\kappa$ a Péclet number defined on the initial size of lamella $s_{0}$.
+Eqs.
+(23), (22) and (8) can be used to predict the spatio-temporal evolution of scalar variance in any mixing front.
+In the following section, we numerically test the accuracy of these theoretical predictions in a periodic, two-dimensional, smooth, and chaotic flow–the random sine flow–for which $D$, $\gamma$ and $\sigma^{2}_{\gamma}$ can be estimated independently.
+Transport is solved for a fixed mean gradient in a periodic domain (Sections  III.3 and III.4) and in the case of a dispersing front in an infinite domain (Section III.5).
 III Comparison to direct numerical simulations
-III.1 Methods
-To numerically test Eq.(34), we first consider advection-diffusion of passive scalar in the presence of a fixed large-scale gradient $\nabla\bar{c}\equiv\bm{g}$.
+To numerically test Eq.(21), we first consider the advection-diffusion of passive scalar in the presence of a fixed large-scale gradient $\nabla\bar{c}\equiv\bm{g}$.
 As done for Eq.(2), the scalar concentration may be split into mean field and fluctuations
@@
 $\bm{g}=g\,\bm{e}_{x},$ (27)
-thus leading to a source term equal to $-gu(y)\bm{e}_{x}$.
-III.2 Velocity field
+thus leading to a source term equal to $gu(y)\bm{e}_{x}$.
+III.1 Velocity field
 We consider a generic smooth chaotic velocity field, the two-dimensional Zeldovich sine flow [34], or random wave flow.
-The sine flow is a continuous transformation operating on a periodic domain $[0,s_{v}]\times[0,s_{v}]$, which is defined as $\bm{u}=(u,v)$, with
+The sine flow is a continuous transformation operating on a periodic domain $0,s_{v}]\times[0,s_{v}]$, which is defined as $\bm{u}=(u,v)$, with
 $v(x,t)$ $=$ $\left\{\begin{array}[]{ll}U\sin(2\pi x/s_{v}+\phi_{j})&\text{ for }jT<t<(j+1/2)T,\\ 0&\text{ for }(j+1/2)T<t<(j+1)T\\ \end{array}\right$.,
@@
 The velocity field is divergence free.
-When $Ku\to 0$ (Eq.
-(3)), the flow changes rapidly over time, causing minimal fluid deformations, while for $Ku\to\infty$ the persistence of the flow strongly affects the deformation of the fluid elements, which tend to align in the main shear directions.
-The limit $Ku\to 0$ is the $\delta$-time correlated flow envisioned in Kraichnan [1], where analytical results about exist (Eq.
-(27)).
-Without loss of generality, we set $T=1$ and $s_{v}=1$ such that the unique free parameter of the flow is $U=Ku$, which controls both the stretching and dispersion statistics of the flow and its persistence.
+When $Ku=UT/s_{v}\to 0$ , the flow changes rapidly over time, causing minimal fluid deformations, while for $Ku\to\infty$ the persistence of the flow strongly affects the deformation of the fluid elements, which tend to align in the main shear directions.
+The limit $Ku\to 0$ is the $\delta$-time correlated flow envisioned in Kraichnan [24], where Eq.
+(11) is valid.
+Without loss of generality, we set $T=1$ and $s_{v}=1$ such that the unique free parameter of the flow is $Ku\equiv U$, which controls both the stretching and dispersion statistics of the flow and its persistence.
 Particle dispersion in random sine flow is given in the long time limit by [14]
 $D\equiv\langle\Delta x^{2}+\Delta y^{2}\rangle/(2T)={U^{2}T}/{16}$ (30)
-where $\Delta x$ and $\Delta y$ are the mean squared displacement caused by a single flow period. In the Kraichnan limit, the mean stretching rate is given by [25, 1]
+where $\Delta x$ and $\Delta y$ are the mean squared displacement caused by a single flow period (see Supp. Mat. for details). In the Kraichnan limit, the mean stretching rate is given by [26, 24]
 $\gamma={\pi^{2}U^{2}T}/({8\,s_{v}^{2}}),$ (31)
-while the variance of stretching rates is equal to the mean, e.g.
-$\sigma_{\gamma}^{2}=\gamma$.
+while the variance of stretching rates is equal to the mean since the flow is two-dimensional, e.g.
+$\sigma_{\gamma}^{2}=\gamma$ (Eq.
+(12) with $d=2$).
 Note that $\gamma\to 0$ as $U\to 0$, e.g.
 in the $\delta$-correlated flow limit.
-In the case of persistent flows ($UT/s_{V}\gg 0$), material lines partially align with shear directions and such that stretching is less efficient and $\gamma<{\pi^{2}U^{2}T}/({8s_{v}^{2}})$.
+In the case of persistent flows ($UT/s_{V}\gg 0$), material lines partially align with shear directions such that stretching is less efficient and $\gamma<{\pi^{2}U^{2}T}/({8s_{v}^{2}})$.
 In addition, the alignement of all material interfaces on the same directions tends to reduce stretching heterogeneity ($\sigma^{2}_{\lambda}<\lambda$).
-More details about this persistent regime can be found in Meunier and Villermaux [36].
-In most random chaotic flows, the elongation of material lines in sine flows approximately follows a log-normal distribution with parameters $\gamma t$ and $\sigma^{2}_{\gamma}t$ that depend on the amplitude $U$, such that Eq.
-(22) is valid.
-III.3 Numerical scheme and convergence
+More details about stretching rates in the persistent regime can be found in Meunier and Villermaux [13].
+III.2 Numerical scheme and convergence
 The advection diffusion equation (7) is solved using spectral decomposition in Fourier space.
@@
 $\tilde{c}_{y}(x,k_{y},t+dt)=\tilde{c}_{y}(x,k_{y},t)\exp(-ik_{y}v(x)\Delta t)$ (35)
-[…]
+followed by a diffusive step (34).
+The only approximation made by the numerical scheme is the operator splitting.
+Thus, the algorithm shows only first-order convergence in $\Delta t$, but with a mean relative error smaller than 0.5% when using a large time step $\Delta t=T/2$.
+The use of GPU makes it extremely efficient, about 2 seconds per time period $T$ for a $20\,000^{2}$ grid with a Nvidia RTX 6000 ADA, which allows us to fully resolve the scalar transport up to Péclet numbers of $10^{8}$ in reasonable time.
 We run simulations for a time large enough to reach stationarity (in a statistical sense).
@@
 This can be explained by the time a scalar fluctuation is transported from large to small scales, at the mean rate $\gamma$.
-Figure 2: a) Time evolution of the scalar variance b) Time convergence of the mean $<\sigma^{2}_{c}>$ over 100 realizations of the random phases for various $U$ at fixed $s_{B}^{\prime}=5.3\cdot 10^{-3}$.
-After this transient, the simulated scalar field reached a statistically steady state, previously reported as strange eigenmode [39] or persistent pattern [21].
-In fact, the asymptotic state of the forced problem is not an eigenmode itself but rather is given by a combination of Floquet modes (Eq.
-(12)) excited by the source term.
+Figure 3: a) Time evolution of the scalar variance b) Time convergence of the mean $\sigma^{2}_{c}>$ over 100 realizations of the random phases for various $U$ at fixed $s_{B}^{\prime}=5.3\cdot 10^{-3}$.
+After this transient regime, the simulated scalar field reached a statistically steady state, previously reported as strange eigenmode [35] or persistent pattern [36].
+In fact, the asymptotic state of the forced problem is not an eigenmode itself but rather is given by a combination of Floquet modes (see Supp.
+Mat.) excited by the source term.
 The spatial structure of this asymptotic mixing state appears different in the presence or absence of large-scale forcing, as shown in Figure 4, with large-scale gradients clearly appearing in the forced case.
@@
 In turn, the persistence of the flow, characterised by the value of $UT/s_{v}$, does not strongly affect the heterogeneity of the concentration field (Fig.
-3), although it clearly impacts the alignment of material lines with respect to the (vertical) velocity field.
+4), although it clearly impacts the alignment of material lines with respect to the (vertical) velocity field.
 Statistical averages are then calculated over time periods of $40\gamma$ and presented below.
-III.4 Spectrum at equilibrium
-Figure 3: Concentration $c^{\prime}$ for various amplitude $U$ for forced mixing $g=1$ (top and scalar decay $g=0$ (bottom).
+III.3 Spectrum at equilibrium
+Figure 4: Concentration $c^{\prime}$ for various amplitude $U$ for forced mixing $g\equiv\nabla\bar{c}=1$ (top and scalar decay $g=0$ (bottom).
 In the scalar decay case, concentrations are rescaled by the standard deviation.
+We used a fixed ratio $\gamma/\kappa=5\cdot 10^{5}$ in all simulations.
 Figure 5: Comparison between power density spectrum obtained via numerical simulation (dots) and Eq.
-(27) (lines) at fixed Péclet number for $U=0.1,0.2$ and $0.8$ (subplots a.1) and a.2)) and at fixed $U=Ku$ and various Péclet numbers (subplots b.1) and b.2)).
+(11) (lines) at fixed $U=0.1$ and various diffusivity (subplots b.1) and b.2)) and at fixed diffusivity $\kappa=10^{-5}$ for increasing flow persistence (subplots a.1) and a.2)).
 The columns show the same spectrum in log-log and log-lin scale.
-We used $\gamma/\chi_{0}=\pi$ to normalise $E_{k}$.
 Note that we normalized $k$ by $1/s_{B}^{\prime}$ in the right column to evidence the universal exponential cutoff.
-We plot the spectrum of $c^{\prime}$ for various flow persistence and Péclet numbers in Figure 4.
-The numerical spectrum clearly shows the $k^{-1}$ scaling at large wavenumber (Figs.
-4a.1 and b.1) and the exponential cutoff at small wavenumbers wavenumber (Figs.
-4a.2 and b.2) .
-We test the accuracy of the Kraichnan model (Eq.
-(27)) combined with the proposed injection mechanism (Eq.(33)) in predicting the numerical spectrum.
+We used the theoretical value of $\chi_{0}$ obtained for a freely dispersing front (Eq.
+(16)) to normalize the spectra.
+Inset of b.1) shows the apparition of spurious frequencies at large flow persistence due to the periodicity of the sine-flow.
+We plot the spectrum of $c^{\prime}$ for weakly persistent flow ($U=0.1$) and various Péclet numbers in Figure 5a.
+The spectrum clearly shows the $k^{-1}$ scaling at small wavenumber (Figs.
+5a.1) and an exponential cutoff at large wavenumbers (Figs.
+5a.2).
 Given the theoretical stretching statistics and dispersivity of the sine flow (Eqs.
@@
 (16)), the scalar variance injection rate is
-${\chi_{0}}=\gamma\,\frac{g^{2}}{\pi}.$ (48)
-[…]
-In the case of persistent flows, the theory accurately captures the spectrum in the small-wavenumber limit but fails to describe the high-wavenumber tails, which significantly depart from the exponential prediction.
-Flow persistence ($U>0$) results in a steeper cutoff than predicted theoretically (Fig.
-4 a.2).
-This may be attributed to the alignment of the material lines in the main shear direction, which tends to limit the heterogeneity of the exponential stretching and push towards a Gaussian cutoff of the spectrum [7].
-We will see later that this is, however, not critical for the prediction of scalar fluctuations.
-III.5 Scalar fluctuations at equilibrium
+$\frac{\chi_{0}}{\gamma}=\frac{(s_{v}\nabla\bar{c})^{2}}{\pi},$ (36)
+in the limit of $\delta$-correlated flow.
+Integrating Eq.
+(36) in Eq.
+(11), we obtain a theoretical PSD that accurately describes the numerical data in the whole range of scales and Péclet numbers for weakly persistent flows ($U\leq 0.1$) (Figs.
+5 b.1 and b.2).
+This confirms that Eq.
+(16) correctly captures the large scale injection mechanism in a freely dispersing front (Eq.
+(19)).
+Note that the prediction of $\chi_{0}/\gamma$ for the sine-flow involves no fitting parameter in contrast to models developed for turbulent flows that use the Batchelor constant as a fitting parameter to link effective to total strain rates [10].
+In the case of persistent flows, the Kraichnan theory accurately captures the spectrum in the small-wavenumber limit (Fig 5b.1) but fails to describe the high-wavenumber tails (Fig 5b.2).
+We observe that flow persistence results in a steeper cutoff of energies than the exponential cutoff predicted by Kraichnan theory.
+Such cutoff may be attributed to the alignment of the material lines in the main shear direction, which tends to limit the spatial heterogeneity of the exponential stretching ($\sigma^{2}_{\gamma}<\gamma$) and push towards a Gaussian cutoff of the spectrum, of Batchelor type [23].
+Note that at high flow persistence, the scalar spectrum tends to a limiting spectrum when plotted against $ks_{B}^{\prime}$, which is well fitted by
+$E_{k}\approx\frac{\chi_{0}}{\gamma k}\exp(-2ks_{B}^{\prime})\exp(-(ks_{B}^{\prime})^{2})$ (37)
+which is in essence Eq.
+(11) augmented by a Gaussian cutoff, à la Batchelor.
+Note that for $U\geq 1$, spurious peaks appear in the spectrum at small wavenumbers (see inset of Fig 5b.1).
+These are caused by the periodicity of the sine flow, which creates an artificial resonance for flow amplitudes larger than the domain size.
+Such peaks would probably disappear in random flows.
+III.4 Scalar fluctuations at equilibrium
 In Fig.
-5, we plot the scalar fluctuation pdf as a function of flow persistence and Péclet number.
+6, we plot the scalar fluctuation pdf for various flow persistence and Péclet numbers.
 Due to the large-scale forcing, the pdf of fluctuation is stationary in time.
 The pdf shows a characteristic Gaussian shape, as previously observed in forced mixing cases [37, 38].
-The Gaussian shape is preserved in all flow persistence and Péclet numbers.
-Note also that the pdf of scalar fluctuation without large-scale forcing ($g=0$) have thicker tails than the Gaussian (Fig.
-5a), which resemble a power law.
-The Gaussianity of scalar fluctuations in mixing fronts is an important feature because it allows one to accurately capture the distribution of fluctuations by the second moments only, that is, the scalar variance.
-Figure 5: Pdf of $c^{\prime}$ as a function of $U$ (Pe=$10^{5}$) and Péclet ($U=0.4$).
+The Gaussian shape is preserved in all flow persistences and Péclet numbers.
+Note also that the pdf of scalar fluctuations without large-scale forcing ($g=0$) have thicker power-law tails than the Gaussian distribution (Fig.
+6a).
+The Gaussianity of scalar fluctuations in mixing fronts is an important feature because it allows one to accurately capture the distribution of fluctuations from the second moment only, that is, the scalar variance.
+Figure 6: Pdf of $c^{\prime}$ as a function of $U$ (with $\kappa=10^{-5}$) and molecular diffusivity (with $U=0.4$).
 The black dotted line correspond to the case without large scale forcing, with the concentration normalized by $\sigma_{c}$.
@@
 In Fig.
-6, we report the value of the scalar variance in several flow and transport conditions and compare it with the theoretical closure proposed herein (Eq.
+7, we compare the value of the simulated scalar variance to the theoretical prediction (Eq.
+(7)) and its approximation with ${F}_{\gamma/\sigma^{2}_{\gamma}}\approx 1$ (Eq.
 (21)).
-A best fit of the closure (38) to the numerical data gives
-$s_{i}\approx 2.7s_{v},$ (49)
-for all flow persistence (Fig.
-6).
-This value is close to the critical domain size that divides globally controlled from locally controlled variance decay rates in sine flows (between 2 and $3s_{v}$  [17, 38]), confirming the validity of the proposed scaling Eq.
-(36).
-Note that the value of $s_{i}$ may differ in other flows than the sine flow.
-Even if Eq.
-(27) relies on the assumption of $\delta$-correlated flows, we found that the prediction for scalar variance is excellent even in the case of correlations.
-This is due to the importance of the universal range $k^{-1}$ in determining the integral (16) when $s_{B}^{\prime}\ll s_{i}$.
-The divergence between curves at large $s_{B}^{\prime}$ (small $(s_{i}/s_{B}^{\prime})^{2}$) is explained by the increasing role of the cutoff shape in the integral, which departs from exponential at finite persistence time (Fig.
-4).
-Figure 6: Dependence of the variance of scalar fluctuations $\sigma^{2}_{c}$ upon modified Batchelor scale $s_{B}^{\prime}$ (Eq.
+The injection scale (Eq.
+(19)) is estimated using Eq.
+(30) and Eq.
+(31) for the dispersion coefficient and the stretching rate respectively, giving
+$s_{i}=2\sqrt{2}\,s_{v}\approx 2.8\,s_{v}.$ (38)
+Note that this value agrees with the critical domain size (between 2 and $3s_{v}$) that divides globally controlled from locally controlled variance decay rates in 2D sine flows [14, 39].
+The predicted variance (Eq.
+21) accurately captures the numerical data over a large range of molecular diffusivity values and flow amplitudes.
+The approximation ${F}_{\gamma/\sigma^{2}_{\gamma}}\approx 1$ is valid even in two-dimensional flows because it is mainly the universal $k^{-1}$ regime at low wavenumbers and the exponential cutoff at high wavenumbers that determines the value of the integral (21), rather than the exact power-law scaling at intermediate wavenumbers.
+While the Kraichnan theory rely on $\delta$-correlated flows, we find that the prediction for scalar variance is excellent even in the case of correlations for $Pe>10^{3}$.
+This is due to the importance of the universal range $k^{-1}$ that emerges both in correlated and uncorrelated flow, which determines the integral (8) when $s_{B}^{\prime}\ll s_{i}$.
+The divergence between curves at large $s_{B}^{\prime}$ (small $Pe$) is explained by the increasing role of the cutoff shape in the integral, which departs from exponential at finite persistence time (see Fig.
+5).
+As shown in Fig.
+7, using the modified spectrum (37) in Eq.
+(8) to account for flow persistence only captures part of the variance drop observed below $Pe=10^{3}$.
+Thus, we conclude that in addition to a change in the spectrum, flow persistence may also push the injection scale $s_{i}$ to smaller values.
+Figure 7: Dependence of the variance of scalar fluctuations $\sigma^{2}_{c}$ upon the Péclet number (Eq.
 (24)) and flow amplitudes ($U$) in the sine flow (colored lines with markers).
-The theoretical approximation Eq.
-(31) is plotted in thick dashed line while the exact numerical integration of Eq.
-(34) is plotted as a thin continuous line.
-In the sine flow, $\chi_{0}/\gamma=1/\pi$ (Eq.
-(48)) The injection scale $s_{i}$ is fixed to $s_{i}=2.7s_{v}$, to match the numerical data.
-It is striking that even at large Péclet numbers, microscale fluctuations remain limited compared to macroscale scalar gradients.
-For example, at $Pe=1000$, scalar fluctuations are expected to be less than 10% of the macroscopic gradient.
+The approximation Eq.
+(21) is plotted in thick dashed line while the exact numerical integration of Kraichnan spectrum (Eq.
+(8)) is plotted as a thin continuous line.
+In the sine flow, $\chi_{0}/\gamma=(s_{v}\nabla\bar{c})^{2}/\pi$ (Eq.
+(36)) and $s_{i}=2\sqrt{2}s_{v}$ (Eq.
+(38)).
+The dashed line represents the numerical integration of the modified Kraichnan spectrum (Eq.
+(37)) observed in the case of persistent flows.
+Note that even at large Péclet numbers, microscale fluctuations remain limited compared to macroscale scalar gradients.
+For example, at $Pe=10^{3}$, scalar fluctuations are expected to be less than 10% of the macroscopic gradient.
 At $Pe=10^{8}$, the fluctuations hardly reach 60% of the macroscopic gradient.
-III.6 Evolution of a mixing front
-Now, we test the capacity of the proposed closure Eq.
-(35) to capture the evolution of a unidirectional mixing front, agitated by a chaotic microscale flow (the sine flow), and that spreads to large scales due to dispersion.
-Since $\bar{c}$ follows the macrodispersion equation with $\overline{\bm{u}}=0$, its solution for an initial sharp step function is
+III.5 Evolution of a mixing front
+We now test the capacity of the proposed closure Eq.
+(21) to capture the evolution in space and time of the scalar variance in a unidirectional mixing front, agitated by a chaotic microscale flow (the sine flow), and that spreads to large scales due to dispersion.
+Since $\bar{c}$ follows the macroscopci dispersion equation with $\overline{\bm{u}}=0$, its solution for an initial sharp step function is
 $\bar{c}=(1+\text{erf}(x/\sqrt{4Dt})/2.$ (39)
-The evolution of the local mean concentration is well captured by the dispersive equation (Fig. 7a)
+The evolution of the local mean concentration is well captured by the dispersive equation (Fig.
+8a).
+The mean concentration is obtained numerically as $\bar{c}(x,t)=\langle c(x,y,t)\rangle_{y}$, where the average is taken on the direction $y$ and on a ensemble of 100 realizations of the random phases $(\phi_{i},\phi_{j})$.
 Figure 8: Direct numerical simulation of the evolution the moments of a scalar front mixed by a chaotic velocity field (continuous line).
@@
 $\sigma^{2}_{c}(x,t)=\frac{\exp(-x^{2}/(2Dt))}{2\pi t}\left|Ei(-4\pi s_{B}^{\prime}/s_{i})\right|.$ (41)
+Numerically, we estimate the variance of scalar fluctuations as $\sigma^{2}_{c}(x,t)=\langle c(x,y,t)^{2}\rangle_{y}-\langle c(x,y,t)\rangle_{y}^{2}$.
 Eq.
 (41) accurately capture the evolution of scalar variance along the front at late time (Fig.
-7b).
+8b).
 At early times ($t\leq 5$), the prediction underestimates the spatial extent of the variance.
@@
 IV Conclusion
-We investigated the statistical behaviour of mixing fronts in smooth chaotic flows in the presence of large-scale scalar gradients.
-We use the spectral description of mixing in time-uncorrelated flows of Kraichnan [1] to propose a new closure model relating the stretching statistics (mean and variance of the stretching rates) to the local variability of solute concentrations (the scalar variance), depending on the strength of large-scale scalar gradient and the Péclet number.
+In this study, we have investigated the statistical behaviour of mixing fronts in chaotic flows, with the objective of describing and predicting the interactions between macroscopic scalar gradients driven by dispersion and microscale mixing controlled by stretching and folding.
+The statistics of concentration fluctuations below the characteristic velocity scale are described the spectral theory of mixing in time-uncorrelated flows of Kraichnan [24] while the macroscopic concentration gradients are described by Fickian dispersion.
+To relate these two descriptions, we identity an injection scale $s_{i}$ defined as the scale at which the characteristic time to destroy variance by dispersion is comparable to the characteristic time to destroy variance by stretching-enhanced mixing.
+This leads to a prediction of the concentration statistics across mixing fronts in chaotic flows.
+The theory captures accurately the results of numerical simulations of mixing fronts in periodic sine flows with no fitting parameters.
 The validity of the approach relies on three key properties of the flow.
-The first is the smoothness of the flow field, which ensures that stretching is a continuous, scale-independent process below the velocity-length scale.
+The first is the smoothness of the flow field, which ensures that stretching is a continuous, scale-independent process below the velocity-length scale and that dispersion is Fickian above the velocity scale.
 The second one is the scale separation induced by chaotic advection, which implies that the dissipation and production scales of the scalar variance are well separated, the Batchelor scale being well below the characteristic velocity length scale.
@@
 We show that this property is, however, not critical, and that a finite amount of flow persistence does not strongly affect the accuracy of the model.
-In addition, recent results [40] show that $\rho$ follows an Orstein-Uhlenbeck process for steady and transient flows, reducing to Eq.
-(22) in the Fickian case.
-The variance of scalar fluctuations created by a chaotic flow in a mean scalar gradient is a universal function of a Péclet number (Fig.
-6), defined as $Pe=\sigma^{2}_{\gamma}s_{i}^{2}/\kappa$, where $\sigma^{2}_{\gamma}$ is the variance of the stretching rates in the chaotic flow (typically of the same order of magnitude of the Lyapunov exponent $\gamma$), $s_{i}$ is a characteristic scale at which scalar variance is produced (typically of the order of the velocity length scale) and $\kappa$ the molecular diffusivity.
-We show that the proposed closure captures direct numerical simulations of advection-diffusion transport in a prototype chaotic flow, the sine flow, and is able to predict the evolution of mean and fluctuation of a solute mixing front.
+In addition, recent results [40] show that $\rho$ follows an Ornstein-Uhlenbeck process for steady and transient flows.
+Analytical results may thus also be within reach for flows with finite persistence.
+One key result of our analysis is that the variance of scalar fluctuations in a freely dispersing mixing front agitated by a smooth chaotic flow is a unique function of a Péclet number comparing dispersive to diffusive fluxes (Eq.
+(24)).
+An explicit expression for the closure is given by Eq.
+(21), and we show that it captures direct numerical simulations of advection-diffusion transport in a prototype chaotic flow, the sine flow, and is able to predict the evolution of mean and fluctuation of a solute mixing front.
 Numerical observation also suggests that local concentration statistics are essentially Gaussian, such that the modelling of the mean and variance of concentrations is sufficient to describe the behaviour of scalar plumes in smooth chaotic flows.
-The microscopic closure (35) is expected to apply in any smooth single-scale chaotic flow of low persistence.
+The closure (21) is expected to apply in any smooth single-scale chaotic flow of moderate persistence.
 This is the case for steady laminar flows, for instance in Stokes flow through 3D porous media, where the velocity fluctuations induce chaotic advection in the 2D plane transverse to mean flow direction [41].
@@
 Thus, flow persistence is of order $Ku\sim 1$.
-In fact, the microscale $s_{m}$ has already been studied in the context of 3D heterogeneous flows at Darcy scale  Kapoor and Gelhar [42].
-However, above pore-scale, advection is in general not chaotic [43], and $s_{m}$ is a non-trivial function of time and space, which strongly dependent on initial conditions.
-In contrast, below the Darcy scale, advection is chaotic and the microscale is a constant, independent of the solute plume geometry.
-Thus, the closure (35) is expected to be relevant in the context of up-scaling transport from pore to Darcy scale in porous media.
-More work is still required to estimate the value of $s_{i}$ in flows other than the sine flow and to confirm its physical origin and universality (Eq.
-(36)).
-Numerical test of the closure in 3D-time dependent chaotic flows is also of direct interest.
-Because it describes the amount of mixing at microscale, the proposed closure opens a new avenue to model non-conservative transport processes, such as reactive processes.
+Because it describes the amount of mixing at microscale in a macroscopically dispersing front, the proposed closure opens a new avenue to upscale non-conservative transport processes, such as reactive processes.
 Extension to bimolecular reactive transport system can be obtained by focusing on the concentration statistics of two reactive species, with fluctuation product $\overline{c_{A}^{\prime}c_{B}^{\prime}}$ rather than a single one $\overline{c^{\prime}c^{\prime}}$ as done here [42].
-This will be the object of a future study.
+This will be the focus of a future study.
+Additional work is also required to validate the proposed expression for $s_{i}$ (Eq.
+(19)) on other flows than the sine flow and to confirm its physical origin and significance in other context.
+Numerical test of the closure in 3D-time dependent chaotic flows is also of direct interest and could be investigated under the same framework.